### v1

# stability of graph pairs involving cycles

Abstract.
A graph pair $(\Gamma,\Sigma)$ is called stable if $\mathrm{Aut}\,(\Gamma)\times\mathrm{Aut}\,(\Sigma)$ is isomorphic to $\mathrm{Aut}\,(\Gamma\times\Sigma)$ and unstable otherwise, where $\Gamma\times\Sigma$ is the direct product of $\Gamma$ and $\Sigma$. A graph is called $R$-thin if distinct vertices have different neighbourhoods. $\Gamma$ and $\Sigma$ are said to be coprime if there is no nontrivial graph $\Delta$ such that $\Gamma\cong\Gamma_{1}\times\Delta$ and $\Sigma\cong\Sigma_{1}\times\Delta$ for some graphs $\Gamma_{1}$ and $\Sigma_{1}$. An unstable graph pair $(\Gamma,\Sigma)$ is called nontrivially unstable if $\Gamma$ and $\Sigma$ are $R$-thin connected coprime graphs and at least one of them is non-bipartite. This paper contributes to the study of the stability of graph pairs with a focus on the case when $\Sigma=C_{n}$ is a cycle. We give two sufficient conditions for $(\Gamma,C_{n})$ to be nontrivially unstable, where $n\neq 4$ and $\Gamma$ is an $R$-thin connected graph. In the case when $\Gamma$ is an $R$-thin connected non-bipartite graph, we obtain the following results: (i) if $(\Gamma,K_{2})$ is unstable, then $(\Gamma,C_{n})$ is unstable for every even integer $n\geq 4$; (ii) if an even integer $n\geq 6$ is compatible with $\Gamma$ in some sense, then $(\Gamma,C_{n})$ is nontrivially unstable if and only if $(\Gamma,K_{2})$ is unstable; (iii) if there is an even integer $n\geq 6$ compatible with $\Gamma$ such that $(\Gamma,C_{n})$ is nontrivially unstable, then $(\Gamma,C_{m})$ is unstable for all even integers $m\geq 6$. We also prove that if $\Gamma$ is an $R$-thin connected graph and $n\geq 3$ is an odd integer compatible with $\Gamma$, then $(\Gamma,C_{n})$ is stable.
Key words and phrases:
stable graph pair, stable graph, expected automorphism, direct product of graphs.
1. Introduction
All graphs considered in this paper are finite and undirected with no loops or parallel edges. As usual, for a graph $\Gamma$, we use $V(\Gamma)$ and $E(\Gamma)$ to denote its vertex set and edge set respectively, and call $|V(\Gamma)|$ the order of $\Gamma$. The edge between two adjacent vertices $u,v$ of $\Gamma$ is denoted by $\{u,v\}$, the neighbourhood of a vertex $u$ in $\Gamma$ is denoted by $N_{\Gamma}(u)$, and the degree of $u$ in $\Gamma$ is defined as $\deg(u)=|N_{\Gamma}(u)|$. Denote by $d(u,v)$ the distance in $\Gamma$ between two vertices $u,v$ of $\Gamma$ and by $\Gamma[S]$ the subgraph of $\Gamma$ induced by a subset $S$ of $V(\Gamma)$. An automorphism of $\Gamma$ is a permutation $\sigma$ of $V(\Gamma)$ such that for any $u,v\in V(\Gamma)$, $\{u,v\}$ is an edge of $\Gamma$ if and only if $\{u^{\sigma},v^{\sigma}\}$ is an edge of $\Gamma$, where for each $w\in V(\Gamma)$, $w^{\sigma}$ is the image of $w$ under $\sigma$. The automorphism group of $\Gamma$, denoted by $\mathrm{Aut}\,(\Gamma)$, is the group of automorphisms of $\Gamma$ under the composition of permutations. A graph $\Gamma$ is called vertex-transitive if for any $u,v\in V(\Gamma)$ there exists an element $\sigma\in\mathrm{Aut}\,(\Gamma)$ such that $u^{\sigma}=v$. We use $K_{n}$ to denote the complete graph with order $n\geq 1$ and $C_{n}$ the cycle with order $n\geq 3$. A graph is trivial if it has only one vertex and nontrivial otherwise.
The direct product [1] of two graphs $\Gamma$ and $\Sigma$, denoted by $\Gamma\times\Sigma$, is the graph with vertex set $V(\Gamma)\times V(\Sigma)$ in which $(u,v)$ and $(x,y)$ are adjacent if and only if $u$ is adjacent to $x$ in $\Gamma$ and $v$ is adjacent to $y$ in $\Sigma$. In particular, $\Gamma\times K_{2}$ is the canonical double cover $D(\Gamma)$ of $\Gamma$. A graph is prime with respect to the direct product if it is nontrivial and cannot be represented as the direct product of two nontrivial graphs. Two graphs $\Gamma$ and $\Sigma$ are said to be coprime with respect to the direct product if there is no nontrivial graph $\Delta$ such that $\Gamma\cong\Gamma^{\prime}\times\Delta$ and $\Sigma\cong\Sigma^{\prime}\times\Delta$ for some graphs $\Gamma^{\prime}$ and $\Sigma^{\prime}$. It is readily seen that for any graphs $\Gamma$ and $\Sigma$, $\mathrm{Aut}\,(\Gamma)\times\mathrm{Aut}\,(\Sigma)$ (direct product of groups) is isomorphic to a subgroup of $\mathrm{Aut}\,(\Gamma\times\Sigma)$. As defined by Qin et al. in [10], a graph pair $(\Gamma,\Sigma)$ is called stable if $\mathrm{Aut}\,(\Gamma)\times\mathrm{Aut}\,(\Sigma)$ is isomorphic to $\mathrm{Aut}\,(\Gamma\times\Sigma)$ and unstable otherwise. This definition generalizes the notion of the stability of a graph [5, 14] in the sense that $(\Gamma,K_{2})$ is stable or unstable if and only if $\Gamma$ is stable or unstable, respectively. The stability of graphs has been studied considerably in the past more than three decades [4, 5, 6, 7, 12, 13, 14]. For example, in [14, Theorems C.1-C.4], Wilson gave four sufficient conditions (see [8] for an amendment to one of them) for a circulant graph to be unstable. In [8, Conjecture 1.3], Qin et al. conjectured that there is no nontrivially unstable circulant of odd order, and in [8, Theorem 1.4] they proved that this is true for circulants of prime order. The stability of the generalized Petersen graphs has been completely determined owing to [14, Theorems P.1-P.2] and [9, Corollary 1.3]. In contrast, the study of the stability of general graph pairs started only recently [10, 11].
A graph $\Gamma$ is called $R$-thin (or vertex-determining) if $N_{\Gamma}(u)\neq N_{\Gamma}(v)$ for any two distinct vertices $u,v$ of $\Gamma$. Graphs which are not $R$-thin are called $R$-thick. An unstable graph is called nontrivially unstable [10, 14] if it is connected, non-bipartite and $R$-thin, and trivially unstable otherwise. In [10, Theorem 1.3], Qin et al. proved that if $(\Gamma,\Sigma)$ is stable, then $\Gamma$ and $\Sigma$ are coprime $R$-thin graphs, and if in addition both $\mathrm{Aut}\,(\Gamma)$ and $\mathrm{Aut}\,(\Sigma)$ are nontrivial groups, then both $\Gamma$ and $\Sigma$ are connected and at least one of them is non-bipartite. Due to this result an unstable graph pair $(\Gamma,\Sigma)$ is called nontrivially unstable [10] if $\Gamma$ and $\Sigma$ are $R$-thin connected coprime graphs and at least one of them is non-bipartite. This definition agrees with the concept of a nontrivially unstable graph introduced in [14] in the sense that $(\Gamma,K_{2})$ is nontrivially unstable if and only if $\Gamma$ is nontrivially unstable.
Let $\Gamma$ and $\Sigma$ be graphs with $V(\Sigma)=\{1,\ldots,n\}$. As in [10, Definition 2.3], we use $\mathrm{P}(\Gamma,\Sigma)$ to denote the set of elements of $\mathrm{Aut}\,(\Gamma\times\Sigma)$ that leave the partition $\{V(\Gamma)\times\{i\}:i\in V(\Sigma)\}$ invariant. Note that $\mathrm{P}(\Gamma,\Sigma)$ is a subgroup of $\mathrm{Aut}\,(\Gamma\times\Sigma)$. An $n$-tuple of permutations $(\alpha_{1},\ldots,\alpha_{n})$ of $V(\Gamma)$ is called [10] a $\Sigma$-automorphism of $\Gamma$ if for all $u,v\in V(\Gamma)$, $\{u,v\}\in E(\Gamma)$ if and only if $\{u^{\alpha_{i}},v^{\alpha_{j}}\}\in E(\Gamma)$ for all $i,j\in V(\Sigma)$ with $\{i,j\}\in E(\Sigma)$. A $\Sigma$-automorphism $(\alpha_{1},\ldots,\alpha_{n})$ of $\Gamma$ is said to be nondiagonal if $\alpha_{i}\neq\alpha_{j}$ for at least one pair of vertices $i,j$ of $\Sigma$. The set of all $\Sigma$-automorphisms of $\Gamma$ with operation defined by $(\alpha_{1},\ldots,\alpha_{n})(\beta_{1},\ldots,\beta_{n})=(\alpha_{1}\beta_{1},\ldots,\alpha_{n}\beta_{n})$ is a group, written $\mathrm{Aut}_{\Sigma}\,(\Gamma)$, which is called [10] the $\Sigma$-automorphism group of $\Gamma$. In particular, $\mathrm{Aut}_{K_{2}}(\Gamma)$ is exactly the two-fold automorphism group of $\Gamma$ (see [2, 3, 4]) and its elements are called the two-fold automorphisms of $\Gamma$.
In [10, Theorem 1.8], Qin et al. proved that, for a connected regular graph $\Gamma$ and a connected vertex-transitive graph $\Sigma$ with coprime degrees, if both $\Gamma$ and $\Sigma$ are $R$-thin and at least one of them is non-bipartite, then $(\Gamma,\Sigma)$ is nontrivially unstable if and only if at least one $\Sigma$-automorphism of $\Gamma$ is nondiagonal. In the same paper they proposed to study the stability of graph pairs $(\Gamma,\Sigma)$ for various special families of graphs $\Gamma$ and/or various special families of graphs $\Sigma$, including the case when $\Sigma$ is $K_{n}$ or $C_{n}$. In [11, Theorem 6], Qin et al. proved that, if $\Gamma$ and $\Sigma$ are regular of coprime degrees and $\Sigma$ is vertex-transitive, then $(\Gamma,\Sigma)$ is nontrivially unstable if and only if $(\Gamma,K_{2})$ is nontrivially unstable. In the same paper they also studied the stability of $(\Gamma,C_{n})$ in the case when $\Gamma$ is regular ([11, Proposition 17]), and they asked under what conditions the pair $(\Gamma,C_{n})$ is nontrivially unstable given that $\Gamma$ is stable and $m\geq 6$ is even ([11, Question 18]).
Motivated by the works above, in this paper we study the stability of $(\Gamma,C_{n})$, where $n\geq 3$ and $\Gamma$ needs not be regular. In Section 2, we prove a few preliminary results on $\Sigma$-automorphisms of $\Gamma$ for a general pair $(\Gamma,\Sigma)$, enriching the theories of $\Sigma$-automorphisms developed in [10]. In particular, we give three necessary conditions for $\Gamma$ to be $R$-thin in terms of nondiagonal $\Sigma$-automorphisms of $\Gamma$ (Lemma 2.3). In Section 3, we study the stability of $(\Gamma,C_{n})$, where $n\neq 4$ and $\Gamma$ is an $R$-thin connected graph which is non-bipartite when $n$ is even. We prove that if $n$ is “compatible” with $\Gamma$ in some sense (Definition 2.5) or $n$ is odd and $N_{\Gamma}(u)\cap N_{\Gamma}(v)$ is not an independent set of $\Gamma$ for any edge $\{u,v\}$ of an auxiliary graph $\Gamma^{*}$ (Definition 2.5) associated with $\Gamma$, then $(\Gamma,C_{n})$ is nontrivially unstable if and only if at least one $C_{n}$-automorphism of $\Gamma$ is nondiagonal (Theorem 3.1). In Section 4, we first prove that if $\Gamma$ is $R$-thin, connected, non-bipartite and unstable, then $(\Gamma,C_{n})$ is unstable for every even integer $n\geq 4$ (Theorem 4.3). We then prove that, for an $R$-thin connected non-bipartite graph $\Gamma$, if an even integer $n\geq 6$ is compatible $\Gamma$, then $(\Gamma,C_{n})$ is nontrivially unstable if and only if $\Gamma$ is unstable, and if there is an even integer $n\geq 6$ compatible with $\Gamma$ such that $(\Gamma,C_{n})$ is nontrivially unstable, then $(\Gamma,C_{m})$ is unstable for all even integers $m\geq 6$ (Theorem 4.5). Finally, we prove that if $\Gamma$ is an $R$-thin connected graph and $n\geq 3$ is an odd integer compatible with $\Gamma$, then $(\Gamma,C_{n})$ is stable (Theorem 4.6).
2. Preliminaries
Let $\Gamma$ and $\Sigma$ be graphs with $V(\Sigma)=\{1,\ldots,n\}$, and let $\phi\in\mathrm{Aut}\,(\Sigma)$. Then
$(\alpha_{1},\ldots,\alpha_{n})^{\phi}=(\alpha_{1^{\phi}},\ldots,\alpha_{n^{\phi}}),\;\text{for $(\alpha_{1},\ldots,\alpha_{n})\in\mathrm{Aut}_{\Sigma}\,(\Gamma)$}$
defines a bijection from $\mathrm{Aut}_{\Sigma}\,(\Gamma)$ to $\mathrm{Aut}_{\Sigma}\,(\Gamma)$.
Lemma 2.1.
Let $\Gamma$ and $\Sigma$ be connected graphs with $V(\Sigma)=\{1,\ldots,n\}$. Then for any $(\alpha_{1},\ldots,\alpha_{n})\in\mathrm{Aut}_{\Sigma}\,(\Gamma)$ and $\phi\in\mathrm{Aut}\,(\Sigma)$ we have $(\alpha_{1},\ldots,\alpha_{n})^{\phi}\in\mathrm{Aut}_{\Sigma}\,(\Gamma)$. Moreover, $\mathrm{Aut}\,(\Sigma)$ is a subgroup of the automorphism group of $\mathrm{Aut}_{\Sigma}\,(\Gamma)$.
Proof.
Since $\phi\in\mathrm{Aut}\,(\Sigma)$, for all $i^{\prime},j^{\prime}\in V(\Gamma)$, there exist $i,j\in V(\Gamma)$ such that $i^{\prime}=i^{\phi}$ and $j^{\prime}=j^{\phi}$. Moreover, $(i,j)$ and $(i^{\prime},j^{\prime})$ determine each other uniquely, and $\{i^{\prime},j^{\prime}\}\in E(\Sigma)$ if and only if $\{i,j\}\in E(\Sigma)$. Since $(\alpha_{1},\ldots,\alpha_{n})\in\mathrm{Aut}_{\Sigma}\,(\Gamma)$, for any $u,v\in V(\Gamma)$, $\{u,v\}\in E(\Gamma)$ if and only if $\{u^{\alpha_{i^{\prime}}},v^{\alpha_{j^{\prime}}}\}\in E(\Gamma)$ for any $i^{\prime},j^{\prime}\in V(\Sigma)$ with $\{i^{\prime},j^{\prime}\}\in E(\Sigma)$. That is, $\{u,v\}\in E(\Gamma)$ if and only if $\{u^{\alpha_{i^{\phi}}},v^{\alpha_{j^{\phi}}}\}\in E(\Gamma)$ for any $i^{\phi},j^{\phi}\in V(\Sigma)$ with $\{i^{\phi},j^{\phi}\}\in E(\Sigma)$. Therefore, $(\alpha_{1},\ldots,\alpha_{n})^{\phi}\in\mathrm{Aut}_{\Sigma}\,(\Gamma)$.
Now for any $\phi\in\mathrm{Aut}\,(\Sigma)$ and $(\alpha_{1},\ldots,\alpha_{n}),(\beta_{1},\ldots,\beta_{n})\in\mathrm{Aut}_{\Sigma}\,(\Gamma)$, set $\gamma_{i}=\alpha_{i}\beta_{i}$ and $i^{\prime}=i^{\phi}$ for $i\in V(\Sigma)$. Since $\phi\in\mathrm{Aut}\,(\Sigma)$, when $i$ ranges over all vertices of $\Sigma$ so does $i^{\prime}$. Hence
$(\alpha_{1},\ldots,\alpha_{n})^{\phi}(\beta_{1},\ldots,\beta_{n})^{\phi}$ $=\,(\alpha_{1^{\phi}},\ldots,\alpha_{n^{\phi}})(\beta_{1^{\phi}},\ldots,\beta_{n^{\phi}})$
$=\,(\alpha_{1^{\prime}}\beta_{1^{\prime}},\ldots,\alpha_{n^{\prime}}\beta_{n^{\prime}})$
$=\,(\gamma_{1^{\prime}},\ldots,\gamma_{n^{\prime}})$
$=\,(\gamma_{1^{\phi}},\ldots,\gamma_{n^{\phi}})$
$=\,(\gamma_{1},\ldots,\gamma_{n})^{\phi}$
$=\,(\alpha_{1}\beta_{1},\ldots,\alpha_{n}\beta_{n})^{\phi}$
$=\,((\alpha_{1},\ldots,\alpha_{n})(\beta_{1},\ldots,\beta_{n}))^{\phi}.$
This means that $\mathrm{Aut}\,(\Sigma)$ is a subgroup of the automorphism group of $\mathrm{Aut}_{\Sigma}\,(\Gamma)$. ∎
In [3, Proposition 3.2], Lauri et al. proved that for any two-fold automorphism $(\alpha_{1},\alpha_{2})$ of a graph $\Gamma$, either both $\alpha_{1}$ and $\alpha_{2}$ are automorphisms of $\Gamma$, or neither of them is. The following is a generalization of this result.
Lemma 2.2.
Let $\Gamma$ and $\Sigma$ be connected graphs with $V(\Sigma)=\{1,\ldots,n\}$. Then for any $(\alpha_{1},\ldots,\alpha_{n})\in\mathrm{Aut}_{\Sigma}\,(\Gamma)$, either $\alpha_{i}\in\mathrm{Aut}\,(\Gamma)$ for all $1\leq i\leq n$ or $a_{i}\notin\mathrm{Aut}\,(\Gamma)$ for all $1\leq i\leq n$.
Proof.
Suppose that $\alpha_{i}\in\mathrm{Aut}\,(\Gamma)$ for some $i$. Then $(\alpha^{-1}_{i},\ldots,\alpha^{-1}_{i})\in\mathrm{Aut}_{\Sigma}\,(\Gamma)$. Since $(\alpha_{1},\ldots,\alpha_{n})\in\mathrm{Aut}_{\Sigma}\,(\Gamma)$ and $\mathrm{Aut}_{\Sigma}\,(\Gamma)$ is a group, we have $(\alpha_{1}\alpha^{-1}_{i},\ldots,\alpha_{i}\alpha^{-1}_{i},\ldots,\alpha_{n}\alpha^{-1}_{i})\in\mathrm{Aut}_{\Sigma}\,(\Gamma)$. Since $\alpha_{i}\alpha^{-1}_{i}={\rm id}$, this shows that we may assume without loss of generality that $\alpha_{i}={\rm id}$ in $(\alpha_{1},\ldots,\alpha_{n})\in\mathrm{Aut}_{\Sigma}\,(\Gamma)$. Then for any $u,v\in V(\Gamma)$ and $j\in N_{\Sigma}(i)$, $\{u,v\}\in E(\Gamma)$ if and only if $\{u^{\alpha_{i}},v^{\alpha_{j}}\}=\{u,v^{\alpha_{j}}\}\in E(\Gamma)$. Moreover, if $\{u,v^{\alpha_{j}}\}\in E(\Gamma)$, then $\{v^{\alpha_{j}},u\}\in E(\Gamma)$ and hence $\{(v^{\alpha_{j}})^{\alpha_{i}},u^{\alpha_{j}}\}=\{v^{\alpha_{j}},u^{\alpha_{j}}\}\in E(\Gamma)$. So we have proved that $\{u,v\}\in E(\Gamma)$ implies $\{u^{\alpha_{j}},v^{\alpha_{j}}\}\in E(\Gamma)$. In other words, $\alpha_{j}$ maps edges of $\Gamma$ to edges of $\Gamma$. Since $\alpha_{j}$ is a permutation of $V(\Gamma)$, we obtain further that $\alpha_{j}$ maps different edges of $\Gamma$ to different edges of $\Gamma$. So $\{u,v\}\notin E(\Gamma)$ implies $\{u^{\alpha_{j}},v^{\alpha_{j}}\}\notin E(\Gamma)$. Therefore, $\alpha_{j}\in\mathrm{Aut}\,(\Gamma)$ for all $j\in N_{\Sigma}(i)$.
If $N_{\Sigma}(j)\subseteq N_{\Sigma}(i)$ for any vertex $j\in N_{\Sigma}(i)$, then $N_{\Sigma}(i)\cup\{i\}=V(\Gamma)$ and so $\alpha_{j}\in\mathrm{Aut}\,(\Gamma)$ for all $j\in N_{\Sigma}(i)$. We now consider the case $V(\Sigma)\setminus N_{\Sigma}(i)\neq\emptyset.$ Define $(\beta_{1},\ldots,\beta_{n})$ such that $\beta_{j}=\alpha_{j}$ for $j\in N_{\Sigma}(i)$ and $\beta_{j}={\rm id}$ for $j\in V(\Sigma)\setminus N_{\Sigma}(i)$. Then for any $u,v\in V(\Gamma)$ and $\{k,l\}\in E(\Sigma)$, $\{u^{\beta_{k}},v^{\beta_{l}}\}$ is equal to $\{u,v\}$, $\{u^{\alpha_{k}},v\}$, $\{u,v^{\alpha_{l}}\}$ or $\{u^{\alpha_{k}},v^{\alpha_{l}}\}$, depending on whether $k\in N_{\Sigma}(i)$ and/or $l\in N_{\Sigma}(i)$. This together with what we proved in the previous paragraph implies that $\{u,v\}\in E(\Gamma)$ if and only if $\{u^{\beta_{k}},v^{\beta_{l}}\}\in E(\Gamma)$. Hence $(\beta_{1},\ldots,\beta_{n})\in\mathrm{Aut}_{\Sigma}\,(\Gamma)$. Since $\mathrm{Aut}_{\Sigma}\,(\Gamma)$ is a group, we then have $(\beta_{1}^{-1},\ldots,\beta_{n}^{-1})\in\mathrm{Aut}_{\Sigma}\,(\Gamma)$ and consequently $(\alpha_{1}\beta_{1}^{-1},\ldots,\alpha_{n}\beta_{n}^{-1})\in\mathrm{Aut}_{\Sigma}\,(\Gamma)$. Note that $\alpha_{j}\beta^{-1}_{j}={\rm id}$ for $j\in N_{\Sigma}(i)$ and $\alpha_{j}\beta^{-1}_{j}=\alpha_{j}$ for $j\in V(\Sigma)\setminus N_{\Sigma}(i)$. Thus, for any $\{u,v\}\in E(\Gamma)$ and any $j\in N_{\Sigma}(i)$ and $l\in N_{\Sigma}(j)$, $\{u,v\}\in E(\Gamma)$ if and only if $\{u^{\alpha_{j}\beta_{j}^{-1}},v^{\alpha_{l}\beta_{l}^{-1}}\}=\{u,v^{\alpha_{l}}\}\in E(\Gamma)$. On the other hand, by $\{u,v^{\alpha_{l}}\}\in E(\Gamma)$ we have $\{v^{\alpha_{l}},u\}\in E(\Gamma)$ and so $\{(v^{\alpha_{l}})^{\alpha_{j}\beta_{j}^{-1}},u^{\alpha_{l}\beta_{l}^{-1}}\}=\{v^{\alpha_{l}},u^{\alpha_{l}}\}\in E(\Gamma)$. Thus $\{u,v\}\in E(\Gamma)$ implies $\{u^{\alpha_{l}},v^{\alpha_{l}}\}\in E(\Gamma)$ for all $l\in N_{\Sigma}(j)$ with $j\in N_{\Sigma}(i)$. That is, $\alpha_{l}$ maps edges of $\Gamma$ to edges of $\Gamma$. Since $\alpha_{l}$ is a permutation of $V(\Gamma)$, it must map different edges of $\Gamma$ to different edges of $\Gamma$. Thus $\{u,v\}\notin E(\Gamma)$ implies $\{u^{\alpha_{l}},v^{\alpha_{l}}\}\notin E(\Gamma)$. Therefore, $\alpha_{l}\in\mathrm{Aut}\,(\Gamma)$ for all $l\in N_{\Sigma}(j)$ with $j\in N_{\Sigma}(i)$. Since $\Sigma$ is a finite connected graph, by repeating the precess a finite number of times we obtain that $\alpha_{j}\in\mathrm{Aut}\,(\Gamma)$ for all $1\leq j\leq n$. ∎
Lemma 2.3.
Let $\Gamma$ and $\Sigma$ be connected graphs with $V(\Sigma)=\{1,\ldots,n\}$. If $\Gamma$ is $R$-thin, then the following statements hold for every nondiagonal $(\alpha_{1},\ldots,\alpha_{n})\in\mathrm{Aut}_{\Sigma}\,(\Gamma)$:
  1. (a)
    $\alpha_{i}\notin\mathrm{Aut}\,(\Gamma)$ for $1\leq i\leq n$;
  2. (b)
    $\alpha_{i}$ for $1\leq i\leq n$ all have the same order;
  3. (c)
    $\alpha_{i}\neq\alpha_{j}$ for every $\{i,j\}\in E(\Sigma)$.
Proof.
Suppose that $\Gamma$ is $R$-thin. Let $(\alpha_{1},\ldots,\alpha_{n})\in\mathrm{Aut}_{\Sigma}\,(\Gamma)$ be an arbitrary nondiagonal $\Sigma$-automorphism of $\Gamma$.
(a) Suppose for a contradiction that $\alpha_{i}\in\mathrm{Aut}\,(\Gamma)$ for at least one $1\leq i\leq n$. Then by Lemma 2.2 we have $\alpha_{i}\in\mathrm{Aut}\,(\Gamma)$ for all $1\leq i\leq n$. Since $\Sigma$ is connected and $(\alpha_{1},\ldots,\alpha_{n})$ is nondiagonal, there exists an edge $\{i,j\}\in E(\Sigma)$ such that $\alpha_{i}\neq\alpha_{j}$ and hence $\gamma:=\alpha_{j}\alpha^{-1}_{i}$ is a non-identity element of $\mathrm{Aut}\,(\Gamma)$. So $\gamma$ moves at least one vertex of $\Gamma$, say, $u$, so that $v:=u^{\gamma}\neq u$. Assume without loss of generality that $i<j$. Then $(\alpha_{1}\alpha_{i}^{-1},\ldots,\alpha_{i}\alpha^{-1}_{i},\ldots,\alpha_{j}\alpha_{i}^{-1},\ldots,\alpha_{n}\alpha_{i}^{-1})\in\mathrm{Aut}_{\Sigma}\,(\Gamma)$ by Lemma 2.1. For each $w\in N_{\Gamma}(u)$, we have $\{w,u^{\gamma}\}=\{w,v\}$. Hence each neighbour $w$ of $u$ in $\Gamma$ is also a neighbour of $v$ in $\Gamma$. So $N_{\Gamma}(u)\subseteq N_{\Gamma}(v)$. By Lemma 2.1 again, we have $((a_{1}a_{i}^{-1})^{-1},\ldots,(a_{i}a_{i}^{-1})^{-1},\ldots,(a_{j}a_{i}^{-1})^{-1},\ldots,(a_{n}a_{i}^{-1})^{-1})\in\mathrm{Aut}_{\Sigma}\,(\Gamma)$. Since $\gamma^{-1}=(\alpha_{j}\alpha_{i}^{-1})^{-1}$, we obtain that $\{w,v^{\gamma^{-1}}\}=\{w,u\}\in E(\Gamma)$ for any $w\in N_{\Gamma}(v)$. In other words, $N_{\Gamma}(v)\subseteq N_{\Gamma}(u)$. Therefore, $N_{\Gamma}(u)=N_{\Gamma}(v)$ and thus $\Gamma$ is not $R$-thin, but this is a contradiction.
(b) Suppose to the contrary that there exist $1\leq i,j\leq n$ such that $\alpha_{i}$ and $\alpha_{j}$ have different orders. Since $\Sigma$ is connected, this assumption implies that there exist $1\leq i,j\leq n$ with $\{i,j\}\in E(\Sigma)$ such that $\alpha_{i}$ and $\alpha_{j}$ have different orders. Without loss of generality we may assume $\{1,n\}\in E(\Sigma)$ and $\alpha_{1},\alpha_{n}$ have orders $p,q$, respectively, for some $p<q$. Since $(\alpha_{1},\ldots,\alpha_{n})\in\mathrm{Aut}_{\Sigma}\,(\Gamma)$, we have $(\alpha_{1}^{p},\alpha_{2}^{p},\ldots,\alpha^{p}_{n})=({\rm id},\alpha_{2}^{p},\ldots,\alpha^{p}_{n})\in\mathrm{Aut}_{\Sigma}\,(\Gamma)$. Moreover, $({\rm id},\alpha_{2}^{p},\ldots,\alpha^{p}_{n})$ is nondiagonal as $\alpha^{p}_{n}\neq{\rm id}$. Since ${\rm id}\in\mathrm{Aut}\,(\Gamma)$, this contradicts what we have proved in part (a).
(c) Suppose that there exists some $\{i,j\}\in E(\Sigma)$ such that $\alpha_{i}=\alpha_{j}$. Since $(\alpha_{1},\ldots,\alpha_{n})\in\mathrm{Aut}_{\Sigma}\,(\Gamma)$ and $\{i,j\}\in E(\Sigma)$, for any $\{u,v\}\in E(\Gamma)$, we have $\{u^{\alpha_{i}},v^{\alpha_{i}}\}=\{u^{\alpha_{i}},v^{\alpha_{j}}\}\in E(\Gamma)$. Hence $\alpha_{i}\in\mathrm{Aut}\,(\Gamma)$. But this contradicts what we have proved in part (a). ∎
Proposition 2.4.
Let $\Gamma$ be an $R$-thin connected graph and $\Sigma$ a connected graph with $V(\Sigma)=\{1,\ldots,n\}$. Let $\Delta$ be a subgraph of $\Sigma$. Then any $(\alpha_{1},\ldots,\alpha_{n})\in\mathrm{Aut}_{\Sigma}\,(\Gamma)$ gives rise to a $\Delta$-automorphism of $\Gamma$, namely $(\alpha_{i})_{i\in V(\Delta)}\in\mathrm{Aut}\,_{\Delta}(\Gamma)$.
Proof.
Since $\Delta$ is a subgraph of $\Sigma$, without loss of generality we may assume $V(\Delta)=\{1,\ldots,m\}$. Then for $i,j\in V(\Delta)$, $\{i,j\}\in E(\Sigma)$ whenever $\{i,j\}\in E(\Delta)$. Since $(\alpha_{1},\ldots,\alpha_{n})\in\mathrm{Aut}_{\Sigma}\,(\Gamma)$, for all $u,v\in V(\Gamma)$ and any $\{i,j\}\in E(\Sigma)$, $\{u,v\}\in E(\Gamma)$ if and only if $\{u^{\alpha_{i}},v^{\alpha_{j}}\}\in E(\Gamma)$. Since $E(\Delta)\subseteq E(\Sigma)$, the same statement holds for $\{i,j\}\in E(\Delta)$ and hence $(\alpha_{i})_{i\in V(\Delta)}\in\mathrm{Aut}\,_{\Delta}(\Gamma)$. ∎
The following definition is essential for establishing our results in the next two sections.
Definition 2.5.
Given a graph $\Gamma$, define $\Gamma^{*}$ to be the graph with vertex set $V(\Gamma)$ in which $u,v\in V(\Gamma)$ are adjacent if and only if $\deg(u)=\deg(v)=2|N_{\Gamma}(u)\cap N_{\Gamma}(v)|$.
For any $u\in V(\Gamma)$, if $u$ is an isolated vertex of $\Gamma^{*}$, define $L(u)=\{0\}$; if $u$ has degree $1$ in $\Gamma^{*}$, define $L(u)=\{0,1\}$; if $u$ has degree at least $2$ in $\Gamma^{*}$, define $L(u)$ to be the set of lengths of the cycles containing $u$ in $\Gamma^{*}$.
We say that an integer $n\geq 3$ is compatible with $\Gamma$ if either $n$ is odd and $n\notin\cap_{u\in V(\Gamma)}L(u)$ or $n$ is even and $n/2\notin\cap_{u\in V(\Gamma)}L(u)$. An integer $n\geq 3$ is incompatible with $\Gamma$ if it is not compatible with $\Gamma$.
Note that if $n$ is compatible with $\Gamma$, then $\Gamma$ and $C_{n}$ must be coprime. Note also that if $\Gamma$ has a vertex of odd degree, then any integer $n\geq 3$ is compatible with $\Gamma$.
3. Sufficient conditions for $(\Gamma,C_{n})$ to be nontrivially unstable
The purpose of this section is to prove the following result. It is in the same spirit as [10, Theorem 1.8] where $\Gamma$ is regular and $\Sigma$ is vertex-transitive. Here $\Gamma$ needs not be regular, but we assume $\Sigma$ is a cycle among other things.
Theorem 3.1.
Let $\Gamma$ be an $R$-thin connected graph and $n\geq 3$ an integer with $n\neq 4$. Suppose that at least one of $\Gamma$ and $C_{n}$ is non-bipartite. Then the following statements hold:
  1. (a)
    if $n$ is compatible with $\Gamma$, then $(\Gamma,C_{n})$ is nontrivially unstable if and only if at least one $C_{n}$-automorphism of $\Gamma$ is nondiagonal;
  2. (b)
    if $n$ is odd and for any two distinct vertices $u,v\in V(\Gamma)$ with $\{u,v\}\in E(\Gamma^{\ast})$, $N_{\Gamma}(u)\cap N_{\Gamma}(v)$ is not an independent set of $\Gamma$, then $(\Gamma,C_{n})$ is nontrivially unstable if and only if at least one $C_{n}$-automorphism of $\Gamma$ is nondiagonal.
We will use the following known results in our proof of Theorem 3.1. Recall that $\mathrm{P}(\Gamma,\Sigma)$ denotes the set of elements of $\mathrm{Aut}\,(\Gamma\times\Sigma)$ which leave the partition $\{V(\Gamma)\times\{i\}:i\in V(\Sigma)\}$ invariant ([10, Definition 2.3]).
Lemma 3.2.
Let $\Gamma$ and $\Sigma$ be graphs.
  1. (a)
    If $\mathrm{Aut}(\Gamma\times\Sigma)=\mathrm{P}(\Gamma,\Sigma)$, then $(\Gamma,\Sigma)$ is unstable if and only if at least one $\Sigma$-automorphism of $\Gamma$ is nondiagonal ([10, Lemma 2.6]).
  2. (b)
    If both $\Gamma$ and $\Sigma$ are connected, $R$-thin and non-bipartite, then $(\Gamma,\Sigma)$ is stable if and only if $\Gamma$ and $\Sigma$ are coprime ([10, Lemma 3.7]).
Denote by $\pi_{1}$ and $\pi_{2}$ the projections from $V(\Gamma\times\Sigma)$ to $V(\Gamma)$ and $V(\Sigma)$, respectively. That is, $(u,i)^{\pi_{1}}=u$ and $(u,i)^{\pi_{2}}=i$ for any $(u,i)\in V(\Gamma\times\Sigma)$. The Boolean square of $\Gamma$, denoted by ${\mathrm{B}}(\Gamma)$, is the graph with vertex set $V(\Gamma)$ and edge set $\{\{u,v\}:u,v\in V(\Gamma),u\neq v,N_{\Gamma}(u)\cap N_{\Gamma}(v)\neq\emptyset\}$. In particular, if $\Gamma$ is the cycle $C_{n}$ with vertices labelled $1,\ldots,n$ consecutively, then ${\mathrm{B}}(C_{n})$ is the graph with vertex set $\{1,\ldots,n\}$ such that each $i$ is adjacent to $i+2$ and $i-2$ modulo $n$. Note that ${\mathrm{B}}(C_{n})$ is a cycle of length $n$ when $n$ is odd, and ${\mathrm{B}}(C_{n})$ is the union of two cycles of length $n/2$ when $n$ is even. Recall that the automorphism group of $C_{n}$ is the dihedral group $D_{2n}=\langle a,b\ |\ a^{n}=b^{2}=e,bab=a^{-1}\rangle$ of order $2n$. For a graph $\Gamma$ and a partition $\mathcal{B}$ of $V(\Gamma)$, the quotient graph $\Gamma_{\mathcal{B}}$ of $\Gamma$ with respect to $\mathcal{B}$ is defined to have vertex set $\mathcal{B}$ such that two blocks $B,C$ of $\mathcal{B}$ are adjacent if and only if there exists at least one edge of $\Gamma$ with one end-vertex in $B$ and the other end-vertex in $C$.
The next technical lemma will be used to prove another lemma (Lemma 3.4) which in turn will be vital to the proof of Theorem 3.1.
Lemma 3.3.
Let $\Gamma$ be an $R$-thin connected graph and $n\geq 3$ an integer with $n\neq 4$. Set $V(C_{n})=\{0,1,\ldots,n-1\}$. Then the following statements hold:
  1. (a)
    for any $u\in V(\Gamma)$, $\{i,j\}\in E({\mathrm{B}}(C_{n}))$ and $\sigma\in\mathrm{Aut}\,(\Gamma\times C_{n})$, if $(u,i)^{\sigma\pi_{2}}\neq(u,j)^{\sigma\pi_{2}}$, then $(u,i)^{\sigma\pi_{1}}=(u,j)^{\sigma\pi_{1}}$;
  2. (b)
    for any $u\in V(\Gamma)$, $\{i,j\}\in E({\mathrm{B}}(C_{n}))$ and $\sigma\in\mathrm{Aut}\,(\Gamma\times C_{n})$, if $(u,i)^{\sigma\pi_{2}}=(u,j)^{\sigma\pi_{2}}$, then $(u,i)^{\sigma\pi_{2}}=(u,k)^{\sigma\pi_{2}}$ where $k\equiv i+2r\pmod{n}$ for any $11\leq r\leq(n/2)-$ 1.
Moreover, if at least one of $\Gamma$ and $C_{n}$ is non-bipartite, then the following statements hold:
  1. (c)
    for any $u\in V(\Gamma)$, $\{i,j\}\in E({\mathrm{B}}(C_{n}))$ and $\sigma\in\mathrm{Aut}\,(\Gamma\times C_{n})$, if $(u,i)^{\sigma\pi_{2}}\neq(u,j)^{\sigma\pi_{2}}$, then $(u,k)^{\sigma\pi_{2}}\neq(u,l)^{\sigma\pi_{2}}$ for any distinct $k,l\in\{1,\ldots,n\}$;
  2. (d)
    for any $\sigma\in\mathrm{Aut}\,(\Gamma\times C_{n})$, if there exist $u\in V(\Gamma)$ and $\{i,j\}\in E({\mathrm{B}}(C_{n}))$ such that $(u,i)^{\sigma\pi_{2}}\neq(u,j)^{\sigma\pi_{2}}$, then $(v,i)^{\sigma\pi_{2}}\neq(v,j)^{\sigma\pi_{2}}$ for every $v\in V(\Gamma)$;
  3. (e)
    if for any $u\in V(\Gamma)$ and $\sigma\in\mathrm{Aut}\,(\Gamma\times C_{n})$ there exists $\{i,j\}\in E({\mathrm{B}}(C_{n}))$ such that $(u,i)^{\sigma\pi_{2}}\neq(u,j)^{\sigma\pi_{2}}$, then for any $v,w\in V(\Gamma)$ joined by an even-length walk in $\Gamma$, any $k\in V(C_{n})$, and any $\delta\in\mathrm{Aut}\,(\Gamma\times C_{n})$, we have $(v,k)^{\delta\pi_{2}}=(w,k)^{\delta\pi_{2}}$;
  4. (f)
    if there exist $u\in V(\Gamma)$ and $\{i,j\}\in E({\mathrm{B}}(C_{n}))$ such that $(u,i)^{\sigma\pi_{2}}\neq(u,j)^{\sigma\pi_{2}}$ for any $\sigma\in\mathrm{Aut}\,(\Gamma\times C_{n})$, then $\mathrm{Aut}(\Gamma\times C_{n})=\mathrm{P}(\Gamma,C_{n})$.
Proof.
In this proof operations on $V(C_{n})=\{0,1,\ldots,n-1\}$ are performed modulo $n$. Note that for any $\{i,j\}\in E({\mathrm{B}}(C_{n}))$, we have $j\equiv i+2\pmod{n}$ or $j\equiv i-2\pmod{n}$, and without loss of generality we may assume $j\equiv i+2\pmod{n}$.
(a) Assume that $u\in V(\Gamma)$, $\{i,j\}\in E({\mathrm{B}}(C_{n}))$ and $\sigma\in\mathrm{Aut}\,(\Gamma\times C_{n})$. Set $(v,k)=(u,i)^{\sigma}$. Taking a vertex $(w,i+1)\in N_{\Gamma\times C_{n}}((u,i))$, we have $(w,i+1)^{\sigma\pi_{2}}=k-1$ or $k+1$. Since $\{(w,i+1),(u,i+2)\}$ is an edge of $\Gamma\times C_{n}$, $(u,i+2)^{\sigma}$ is adjacent to $(w,i+1)^{\sigma}$. Hence $(u,i+2)^{\sigma\pi_{2}}=k-2$, $k+2$ or $k$. Thus, either $(u,i+2)^{\sigma\pi_{2}}\in\{k+2,k-2\}$ or $(u,j)^{\sigma\pi_{2}}=k$. Note that $(u,i)^{\sigma\pi_{2}}\neq(u,j)^{\sigma\pi_{2}}$ by our assumption. Consider first the case where $(u,j)^{\sigma\pi_{2}}\equiv k+2\pmod{n}$. We aim to prove $(u,i)^{\sigma\pi_{1}}=(u,j)^{\sigma\pi_{1}}$. Suppose otherwise. Then $(u,j)^{\sigma}=(w,k+2)$ for some $w\in V(\Gamma)$. We have
$|N_{\Gamma\times C_{n}}((u,i)^{\sigma})\cap N_{\Gamma\times C_{n}}((u,j)^{\sigma})|$ $=\,|N_{\Gamma\times C_{n}}((v,k))\cap N_{\Gamma\times C_{n}}((w,k+2))|$
$=\,|N_{\Gamma}(v)\cap N_{\Gamma}(w)|\cdot|N_{C_{n}}(k)\cap N_{C_{n}}(k+2)|$
$=\,|N_{\Gamma}(v)\cap N_{\Gamma}(w)|.$
On the other hand,
$|N_{\Gamma\times C_{n}}((u,i)^{\sigma})\cap N_{\Gamma\times C_{n}}((u,j)^{\sigma})|$ $=\,|N_{\Gamma\times C_{n}}((u,i))\cap N_{\Gamma\times C_{n}}((u,j+2))|$
$=\,|N_{\Gamma}(u)\cap N_{\Gamma}(u)|\cdot|N_{C_{n}}(i)\cap N_{C_{n}}(j+2)|$
$=\,|N_{\Gamma}(u)\cap N_{\Gamma}(u)|$
$=\,|N_{\Gamma}(u)|.$
Thus, $|N_{\Gamma}(u)|=|N_{\Gamma}(v)|\cap|N_{\Gamma}(w)|$, but this contradicts the assumption that $\Gamma$ is $R$-thin. This contradiction shows that we have $(u,i)^{\sigma\pi_{1}}=(u,j)^{\sigma\pi_{1}}$ when $(u,j)^{\sigma\pi_{2}}\equiv k+2\pmod{n}$. Similarly, we can prove $(u,i)^{\sigma\pi_{1}}=(u,j)^{\sigma\pi_{1}}$ when $(u,i+2)^{\sigma\pi_{2}}\equiv k-2\pmod{n}$.
(b) Assume that $u\in V(\Gamma)$, $\{i,j\}\in E({\mathrm{B}}(C_{n}))$ and $\sigma\in\mathrm{Aut}\,(\Gamma\times C_{n})$. We proceed by induction on $r$. If $r=1$, then $k=i+2$. Thus $k=j$ and we obtain the result by the assumption. Suppose that, for some $m\geq 1$, the result is true for every $r$ between $1$ and $m-1$. That is, if $(u,i)^{\sigma\pi_{2}}=(u,j)^{\sigma\pi_{2}}$, then $(u,i)^{\sigma\pi_{2}}=(u,i+2r)^{\sigma\pi_{2}}$ for $11\leq r\leq m-$ 1. We aim to prove that the result is also true when $r=m$. By our hypothesis, it suffices to prove $(u,i+2m-2)^{\sigma\pi_{2}}=(u,i+2m)^{\sigma\pi_{2}}$. Suppose for a contradiction that $(u,i+2m-2)^{\sigma\pi_{2}}\neq(u,i+2m)^{\sigma\pi_{2}}$. Then by part (a) we have $(u,i+2m-2)^{\sigma\pi_{1}}=(u,i+2m)^{\sigma\pi_{1}}$. Since $(u,i+2m-2)^{\sigma\pi_{2}}\neq(u,i+2m)^{\sigma\pi_{2}}$, either $(u,i+2m)^{\sigma\pi_{2}}=(u,i+2m-2)^{\sigma\pi_{2}}+2$ or $(u,i+2m)^{\sigma\pi_{2}}=(u,i+2m-2)^{\sigma\pi_{2}}-2$. Consider the former case first. In this case we have $(u,i+2m-2)^{\sigma\pi_{2}}=l$ and $(u,i+2m)^{\sigma\pi_{2}}\equiv l+2\pmod{n}$ for some $l\in V(C_{n})$. So $(u,i+2m-4)^{\sigma\pi_{2}}=(u,i+2m-2)^{\sigma\pi_{2}}=l$. Thus there exists a vertex $(v,i+2m-3)\in N_{\Gamma\times C_{n}}((u,i+2m-4))\cap N_{\Gamma\times C_{n}}((u,i+2m-2))$ such that $(v,i+2m-3)^{\sigma\pi_{2}}\equiv l+1\pmod{n}$ and $(v,i+2m-3)^{\sigma}\in N_{\Gamma\times C_{n}}((u,i+2m-4)^{\sigma})\cap N_{\Gamma\times C_{n}}((u,i+2m-2)^{\sigma})$. So $(v,i+2m-3)^{\sigma}\in N_{\Gamma\times C_{n}}((u,i+2m)^{\sigma})$, but this contradicts the fact that $(v,i+2m-3)$ is not adjacent to $(u,i+2m)$. Similarly, one can derive a contradiction in the case when $(u,i+2m)^{\sigma\pi_{2}}=(u,i+2m-2)^{\sigma\pi_{2}}-2$. Therefore, we must have $(u,i+2m-2)^{\sigma\pi_{2}}=(u,i+2m)^{\sigma\pi_{2}}$ and hence the result is true when $r=m$. This completes the proof by mathematical induction.
In the rest of the proof we assume that at least one of $\Gamma$ and $C_{n}$ is non-bipartite. Equivalently, $n$ is odd when $\Gamma$ is bipartite.
(c) Suppose that $u\in V(\Gamma)$, $\{i,j\}\in E({\mathrm{B}}(C_{n}))$ and $\sigma\in\mathrm{Aut}\,(\Gamma\times C_{n})$ such that $(u,i)^{\sigma\pi_{2}}\neq(u,j)^{\sigma\pi_{2}}$, where $j\equiv i+2\pmod{n}$. Set $(u,i)^{\sigma\pi_{2}}=i_{1}$ and assume without loss of generality that $(u,i+2)^{\sigma\pi_{2}}\equiv i_{1}+2\pmod{n}$. Since both $\Gamma$ and $C_{n}$ are connected and at least one of them is non-bipartite, $\Gamma\times C_{n}$ is connected. So for any $(u,i)\in V(\Gamma\times C_{n})$, there exists a walk from $(u,i)$ to $(u,i+1)$ with vertices in $V(\Gamma)\times\{i,i+1\}$, say, $(u,i)=(u_{0},i),(u_{1},i+1),(u_{2},i),\ldots,(u_{2r},i),(u_{2r+1},i+1)=(u,i+1)$, for some $r$ between 00 and $|V(\Gamma)|-1$. Since $(u,i)^{\sigma\pi_{2}}=i_{1}$ and $(u,i+2)^{\sigma\pi_{2}}=i_{1}+2\pmod{n}$, for each $v\in N_{\Gamma}(u)$ we have $(v,i+1)^{\sigma\pi_{2}}\equiv i_{1}+1\pmod{n}$ and $(v,i-1)^{\sigma\pi_{2}}\equiv i_{1}-1\pmod{n}$. Combining this with $u_{1}\in N_{\Gamma}(u)$, we obtain that $(u_{1},i+1)^{\sigma\pi_{2}}\equiv i_{1}+1\pmod{n}$. By part (b), we have $(u_{1},i+1)\neq(u_{1},i+1+2r)$ for any integer $r$ between 00 and $n/2$. Thus for each $w\in N_{\Gamma}(u_{1})$ we have $(w,i)^{\sigma\pi_{2}}=i_{1}$ and $(w,i+2)^{\sigma\pi_{2}}\equiv i_{1}+2\pmod{n}$. Hence $(u_{2},i)^{\sigma\pi_{2}}=i_{1}$. Applying successively the same argument to pairs of every other vertices in the walk above, we obtain that $(u_{2t+1},i+1)^{\sigma\pi_{2}}\equiv i_{1}+1\pmod{n}$ for $0\leq t\leq r$ and $(u_{2t},i)^{\sigma\pi_{2}}=i_{1}$ for $1\leq t\leq r$. Thus $(u,i+1)^{\sigma\pi_{2}}\equiv i_{1}+1\pmod{n}$. Similarly, we can prove that for any $0\leq t\leq n$, $(u,i+t)^{\sigma\pi_{2}}\equiv i_{1}+t\pmod{n}$ when $(u,i)^{\sigma\pi_{2}}=i_{1}$ and $(u,i+2)^{\sigma\pi_{2}}\equiv i_{1}+2\pmod{n}$. Hence $(u,k)^{\sigma\pi_{2}}\neq(u,l)^{\sigma\pi_{2}}$ for any distinct $k,l\in\{1,\ldots,n\}$.
(d) Let $\sigma\in\mathrm{Aut}\,(\Gamma\times C_{n})$. Assume that there exist $u\in V(\Gamma)$ and $\{i,j\}\in E({\mathrm{B}}(C_{n}))$ such that $(u,i)^{\sigma\pi_{2}}\neq(u,j)^{\sigma\pi_{2}}$. We aim to prove $(v,i)^{\sigma\pi_{2}}\neq(v,j)^{\sigma\pi_{2}}$ for all $v\in V(\Gamma)$. Since this inequality holds when $v=u$, we assume $v\neq u$ in the sequel. Since $(u,i)^{\sigma\pi_{2}}\neq(u,j)^{\sigma\pi_{2}}$, by part (c), we have $(u,k)^{\sigma\pi_{2}}\neq(u,l)^{\sigma\pi_{2}}$ for any distinct $k,l\in\{0,1,\ldots,n-1\}$. Thus, if $u$ and $v$ are adjacent in $\Gamma$, then $(u,i-1)^{\sigma\pi_{2}}$, $(u,i+1)^{\sigma\pi_{2}}$ and $(u,i+3)^{\sigma\pi_{2}}$ are pairwise distinct, and hence $(v,i)^{\sigma\pi_{2}}\neq(v,i+2)^{\sigma\pi_{2}}$. By part (c) again, we have $(v,k)^{\sigma\pi_{2}}\neq(v,l)^{\sigma\pi_{2}}$ for any distinct $k,l\in\{0,1,\ldots,n-1\}$, and so $(v,i)^{\sigma\pi_{2}}\neq(v,j)^{\sigma\pi_{2}}$. If $u$ and $v$ are not adjacent in $\Gamma$, then there is a path in $\Gamma$ from $u$ to $v$ with length at least $2$, say, $u=u_{0},u_{1},\ldots,u_{p}=v$. As shown above, we have $(u_{1},i)^{\sigma\pi_{2}}\neq(u_{1},i+2)^{\sigma\pi_{2}}$. Similarly to the proof in the case when $u$ and $v$ are adjacent, we can prove $(u_{2},i)^{\sigma\pi_{2}}\neq(u_{2},i+2)^{\sigma\pi_{2}}$. Applying successively what we have proved to other vertices in the path $u=u_{0},u_{1},\ldots,u_{p}=v$, we obtain $(u_{k},i)^{\sigma\pi_{2}}\neq(u_{k},i+2)^{\sigma\pi_{2}}$ for $0\leq k\leq p$. In particular, we have $(v,i)^{\sigma\pi_{2}}\neq(v,i+2)^{\sigma\pi_{2}}$. By part (c), we then obtain $(v,i)^{\sigma\pi_{2}}\neq(v,j)^{\sigma\pi_{2}}$ as desired.
(e) Assume that for any $u\in V(\Gamma)$ and $\sigma\in\mathrm{Aut}\,(\Gamma\times C_{n})$ there exists $\{i,j\}\in E({\mathrm{B}}(C_{n}))$ such that $(u,i)^{\sigma\pi_{2}}\neq(u,j)^{\sigma\pi_{2}}$. Let $v,w\in V(\Gamma)$ be such that $d(v,w)=2r$ is even. Consider the case $r=1$ first. In this case we have $N_{\Gamma}(v)\cap N_{\Gamma}(w)\neq\emptyset$ and so we may take $x\in N_{\Gamma}(v)\cap N_{\Gamma}(w)$. For any $k\in V(C_{n})$, set $W=\{(x,k-1),(x,k+1)\}$. Then $W\subseteq N_{\Gamma\times C_{n}}((v,k))\cap N_{\Gamma\times C_{n}}((w,k))$. Since $\sigma\in\mathrm{Aut}\,(\Gamma\times C_{n})$, we have $W^{\sigma}=\{(x,k-1)^{\sigma},(x,k+1)^{\sigma}\}\subseteq N_{\Gamma\times C_{n}}((v,k)^{\sigma})\cap N_{\Gamma\times C_{n}}((w,k)^{\sigma})$. Since $(u,i)^{\sigma\pi_{2}}\neq(u,j)^{\sigma\pi_{2}}$, by part (d), we have $(x,i)^{\sigma\pi_{2}}\neq(x,j)^{\sigma\pi_{2}}$ for any $x\in V(\Gamma)$. Moreover, by part (c), we have $(x,k-1)^{\sigma\pi_{2}}\neq(x,k+1)^{\sigma\pi_{2}}$. Combining this with part (a), we obtain $(x,k-1)^{\sigma\pi_{1}}=(x,k+1)^{\sigma\pi_{1}}$. Note that $(v,k)^{\sigma}$ and $(w,k)^{\sigma}$ are in $N_{\Gamma\times C_{n}}((x,k-1)^{\sigma})\cap N_{\Gamma\times C_{n}}((x,k+1)^{\sigma})$ and $(x,k-1)^{\sigma\pi_{1}}=(x,k+1)^{\sigma\pi_{1}}$. Hence either $(v,k)^{\sigma\pi_{2}}=(w,k)^{\sigma\pi_{2}}$ or $C_{n}=C_{4}$. Since $n\neq 4$ by our assumption, we have $(v,k)^{\sigma\pi_{2}}=(w,k)^{\sigma\pi_{2}}$ as required. Now assume $r\geq 2$. Since $d(v,w)=2r$, there is a path in $\Gamma$ from $v$ to $w$ with length $2r$. Applying what we have proved to pairs of every other vertices in this path, we can obtain the desired result.
(f) Suppose that there exist $u\in V(\Gamma)$ and $\{i,j\}\in E({\mathrm{B}}(C_{n}))$ such that $(u,i)^{\sigma\pi_{2}}\neq(u,j)^{\sigma\pi_{2}}$ for any $\sigma\in\mathrm{Aut}\,(\Gamma\times C_{n})$. We may assume $j\equiv i+2\pmod{n}$ without loss of generality. By part (d), we have $(v,i)^{\sigma\pi_{2}}\neq(v,j)^{\sigma\pi_{2}}$ for any $v\in V(\Gamma)$. Moreover, by part (c), we have $(v,k)^{\sigma\pi_{2}}\neq(v,l)^{\sigma\pi_{2}}$ for any distinct $k,l\in\{0,1,\cdots,n-1\}$.
Consider first the case where $n$ is odd. We aim to prove $(v,k)^{\sigma\pi_{2}}=(w,k)^{\sigma\pi_{2}}$ for any distinct $v,w\in V(\Gamma)$, any $k\in V(C_{n})$, and any $\sigma\in\mathrm{Aut}(\Gamma\times C_{n})$. If there is a walk from $v$ to $w$ in $\Gamma$ with even length, then we obtain the result by part (e). Assume there is no such a walk in the sequel. Consider the case where $v$ and $w$ are adjacent first. In this case, for any $i\in V(C_{n})$, we can construct a cycle $C_{v,w}$ of length $2n$ in $\Gamma\times C_{n}$ containing $(v,i)$, namely
$C_{v,w}:(v,i),(w,i+1),(v,i+2),(w,i+3)\cdots,(v,i-1),(w,i),(v,i+1),\cdots,(w,i-1),(v,i).$
Since $(v,i)^{\sigma\pi_{2}}\neq(v,i+2)^{\sigma\pi_{2}}$ for any $\sigma\in\mathrm{Aut}\,(\Gamma\times C_{n})$, by part (a) we have $(v,i)^{\sigma\pi_{1}}=(v,i+2)^{\sigma\pi_{1}}$. Since $C_{v,w}$ is a cycle in $\Gamma\times C_{n}$ and $(u,r)$ and $(w,r)$ are antipodal in $C_{v,w}$ for $10\leq r\leq n-$ 1, $C_{v,w}^{\sigma}$ is a cycle in $\Gamma\times C_{n}$ and $(v,i)^{\sigma\pi_{2}}$ and $(w,i)^{\sigma\pi_{2}}$ are antipodal in $C_{v,w}^{\sigma}$. Therefore, $(v,i)^{\sigma\pi_{2}}=(w,i)^{\sigma\pi_{2}}$. Now consider the case where $v$ and $w$ are not adjacent. Since $\Gamma$ is connected, there exists a walk from $v$ to $w$ of odd length in $\Gamma$, say, $v=v_{0},v_{1},\ldots,v_{2r+1}=w$, for some $r\geq 1$. Then for any $\sigma\in\mathrm{Aut}\,(\Gamma\times C_{n})$, as shown above, we have $(v_{1},k)^{\sigma\pi_{2}}=(v_{0},k)^{\sigma\pi_{2}}$ for any $k\in V(C_{n})$. Since there is a walk from $v_{1}$ to $v_{2r+1}$ with even length, we have $(v_{1},k)^{\sigma\pi_{2}}=(v_{2r+1},k)^{\sigma\pi_{2}}$ for any $k\in V(C_{n})$. Thus, $(v_{0},k)^{\sigma\pi_{2}}=(v_{2r+1},k)^{\sigma\pi_{2}}$, that is, $(v,k)^{\sigma\pi_{2}}=(w,k)^{\sigma\pi_{2}}$, for any $k\in V(C_{n})$, which implies that $\sigma\in\mathrm{P}(\Gamma,C_{n})$ for any $\sigma\in\mathrm{Aut}(\Gamma\times C_{n})$. Hence $\mathrm{Aut}(\Gamma\times C_{n})=\mathrm{P}(\Gamma,C_{n})$ when $n$ is odd.
Now consider the case where $n\geq 4$ is even. Since by our assumption at least one of $\Gamma$ and $C_{n}$ is non-bipartite, in this case $\Gamma$ must be non-bipartite. Set $\Delta=\Gamma\times C_{n}$ and =$\mathcal{B}=\{V(\Gamma)\times\{j\}:j\in V(C_{n})\}$. Then the quotient graph $\Delta_{\mathcal{B}}$ of $\Delta$ with respect to the partition $\mathcal{B}$ of $V(\Delta)$ is isomorphic to $C_{n}$. Since $n$ is even, $C_{n}$ is bipartite and hence $\Gamma\times C_{n}$ is bipartite. Moreover, since $\Gamma\times C_{n}$ is connected, any two vertices $(v,i),(w,i)\in V(\Gamma\times C_{n})$ are joined by a path of even length in $\Gamma\times C_{n}$, which implies that there is a walk of even length from $v$ to $w$ in $\Gamma$. Thus, by part (e), we obtain that any $\sigma\in\mathrm{Aut}(\Gamma\times C_{n})$ satisfies $(v,k)^{\sigma\pi_{2}}=(w,k)^{\sigma\pi_{2}}$ for any $v,w\in V(\Gamma)$ and $k\in V(C_{n})$, and hence any $\sigma\in\mathrm{Aut}(\Gamma\times C_{n})$ belongs to $\mathrm{P}(\Gamma,C_{n})$. Therefore, $\mathrm{Aut}\,(\Gamma\times C_{n})\subseteq\mathrm{P}(\Gamma,C_{n})$ and consequently $\mathrm{Aut}\,(\Gamma\times C_{n})=\mathrm{P}(\Gamma,C_{n})$ when $n$ is even. ∎
The following lemma plays a key role in our proof of Theorem 3.1. It will also be used to prove two results (Theorems 4.5 and 4.6) in the next section.
Lemma 3.4.
Let $\Gamma$ be an $R$-thin connected graph and $n\geq 3$ an integer with $n\neq 4$. Suppose that at least one of $\Gamma$ and $C_{n}$ is non-bipartite. Then the following statements hold:
  1. (a)
    if $n$ is compatible with $\Gamma$, then $\mathrm{Aut}(\Gamma\times C_{n})=\mathrm{P}(\Gamma,C_{n})$;
  2. (b)
    if $n$ is odd and for any $u,v\in V(\Gamma)$ with $\{u,v\}\in E(\Gamma^{*})$, $N_{\Gamma}(u)\cap N_{\Gamma}(v)$ is not an independent set of $\Gamma$, then $\mathrm{Aut}(\Gamma\times C_{n})=\mathrm{P}(\Gamma,C_{n})$.
Proof.
Set $V(C_{n})=\{0,1,\ldots,n-1\}$, in which operations are taken modulo $n$. Recall that for any $\{i,j\}\in E({\mathrm{B}}(C_{n}))$, $j\equiv i+2\pmod{n}$ or $j\equiv i-2\pmod{n}$. Without loss of generality we may assume $j\equiv i+2\pmod{n}$ whenever $\{i,j\}\in E({\mathrm{B}}(C_{n}))$. By Lemma 3.3(f), we only need to prove the existence of $u\in V(\Gamma)$ and $\{i,j\}\in E({\mathrm{B}}(C_{n}))$ such that $(u,i)^{\sigma\pi_{2}}\neq(u,j)^{\sigma\pi_{2}}$ for any $\sigma\in\mathrm{Aut}(\Gamma\times C_{n})$.
(a) We only prove the result when $n$ is even as the proof for odd $n$ is similar. Since $n$ is even and $n$ is compatible with $\Gamma$, we may take a vertex $u\in V(\Gamma)$ such that $n/2\notin L(u)$. Suppose for a contradiction that there exist $(v,i)\in V(\Gamma\times C_{n})$ and $\sigma\in\mathrm{Aut}\,(\Gamma\times C_{n})$ such that $(v,i)^{\sigma\pi_{2}}=(v,i+2)^{\sigma\pi_{2}}$. Then by Lemma 3.3(d) we have $(w,i)^{\sigma\pi_{2}}=(w,i+2)^{\sigma\pi_{2}}$ for each $w\in V(\Gamma)$. Set $u=(w,i)^{\sigma\pi_{1}}$ and $m=(n/2)-1$. By Lemma 3.3(b), for any $r$ between 00 and $m$, we have $(w,i)^{\sigma\pi_{2}}=(w,k)^{\sigma\pi_{2}}$, where $k\equiv i+2r\pmod{n}$. Thus there are $n/2$ distinct vertices $u=u_{0},u_{1},\ldots,u_{m}\in V(\Gamma)$ and one vertex $l\in V(C_{n})$ such that $(w,k)^{\sigma}=(u_{r},l)$ where $k\equiv i+2r\pmod{n}$ for $0\leq r\leq m$. Since $\sigma\in\mathrm{Aut}\,(\Gamma\times C_{n})$, it follows that $\deg((w,k))=\deg((u_{r},l))$. Moreover, since $C_{n}$ is a regular graph with degree $2$, we obtain further that $\deg(w)=\deg(u_{r})$ for $0\leq r\leq m$. Note that for $0\leq r\leq m$ we have
$|N_{\Gamma\times C_{n}}((w,i+2r)^{\sigma})\cap N_{\Gamma\times C_{n}}((w,i+2r+2)^{\sigma})|$ $=\,|N_{\Gamma\times C_{n}}((u_{r},l))\cap N_{\Gamma\times C_{n}}((u_{r+1},l))|$
$=\,|N_{\Gamma}(u_{r})\cap N_{\Gamma}(u_{r+1})|\cdot|N_{C_{n}}(l)|$
= $=\,2|N_{\Gamma}(u_{r})\cap N_{\Gamma}(u_{r+1})|.$
On the other hand,
$|N_{\Gamma\times C_{n}}((w,i+2r)^{\sigma})\cap N_{\Gamma\times C_{n}}((w,i+2r+2)^{\sigma})|$ $=\,|N_{\Gamma\times C_{n}}((w,i+2r))\cap N_{\Gamma\times C_{n}}((w,i+2r+2))|$
$=\,|N_{\Gamma}(w)\cap N_{\Gamma}(w)|\cdot|N_{C_{n}}(i+2r)\cap N_{C_{n}}(i+2r+2)|$
$=\,|N_{\Gamma}(w)\cap N_{\Gamma}(w)|$
$=\,|N_{\Gamma}(w)|.$
Thus $N_{\Gamma}(w)=2|N_{\Gamma}(u_{r})\cap N_{\Gamma}(u_{r+1})|$ for $0\leq r\leq m$. So there is a cycle in $\Gamma^{*}$ with length $n/2$ which contains $u$, but this is a contradiction. Therefore, we have $(w,i)^{\sigma\pi_{2}}\neq(w,j)^{\sigma\pi_{2}}$ for any $\{i,j\}\in E({\mathrm{B}}(C_{n}))$. Since this holds for any $\sigma\in\mathrm{Aut}\,(\Gamma\times C_{n})$, by Lemma 3.3(e) we obtain that $(v,i)^{\sigma\pi_{2}}\neq(v,i+2)^{\sigma\pi_{2}}$ for any $(v,i)\in V(\Gamma\times C_{n})$ and $\sigma\in\mathrm{Aut}\,(\Gamma\times C_{n})$. This together with Lemma 3.3(f) yields $\mathrm{Aut}(\Gamma\times C_{n})=\mathrm{P}(\Gamma,C_{n})$.
(b) Suppose, for a contradiction, that there exist $(u,i)\in V(\Gamma\times C_{n})$ and $\sigma\in\mathrm{Aut}(\Gamma\times C_{n})$ such that $(u,i)^{\sigma\pi_{2}}=(u,i+2)^{\sigma\pi_{2}}$. Since $n$ is odd, by Lemma 3.3(b), we have $(u,i)^{\sigma\pi_{2}}=(u,k)^{\sigma\pi_{2}}$ for any $k\in V(C_{n})$. So there are vertices $(v,k),(v,k+2)\in V(\Gamma\times C_{n})$ such that $(u,i)^{\sigma}=(v,k)$ and $(u,i+2)^{\sigma}=(w,k)$. Note that $\deg(u)=\deg(v)=\deg(w)$ and $\deg(u)=2|N_{\Gamma}(v)\cap N_{\Gamma}(w)|$. By our assumption, $N_{\Gamma}(v)\cap N_{\Gamma}(w)$ contains two adjacent vertices, say, $x$ and $y$. By parts (b) and (d) of Lemma 3.3, we obtain that $(v,l)^{\sigma\pi_{2}}=(v,l+2)^{\sigma\pi_{2}}$ and $(w,l)^{\sigma\pi_{2}}=(w,l+2)^{\sigma\pi_{2}}$ for any $l\in V(C_{n})$. Moreover, since $n$ is odd, we have $(v,l)^{\sigma\pi_{2}}=(v,l+2r)^{\sigma\pi_{2}}$ and $(w,l)^{\sigma\pi_{2}}=(w,l+2r)^{\sigma\pi_{2}}$ for $10\leq r\leq(n/2)-$ 1. Let $k_{1}=(v,l)^{\sigma\pi_{2}}$ and $k_{2}=(w,l)^{\sigma\pi_{2}}$. Since $N_{\Gamma}(v)\cap N_{\Gamma}(w)\neq\emptyset$, either $k_{1}=k_{2}$ or $k_{1}\in\{k_{2}+2,k_{2}-2\}$.
By parts (b) and (d) of Lemma 3.3 again, we also obtain that $(x,l)^{\sigma\pi_{2}}=(x,l+2r)^{\sigma\pi_{2}}$ and $(y,l)^{\sigma\pi_{2}}=(y,l+2r)^{\sigma\pi_{2}}$ for any $l\in V(C_{n})$ and $10\leq r\leq(n/2)-$ 1. Let $k_{3}=(x,l)^{\sigma\pi_{2}}$ and $k_{4}=(y,l)^{\sigma\pi_{2}}$.
Since $x$ is adjacent to $y$, $(x,l)$ is adjacent to $(y,l+1)$ and $(y,l-1)$. Hence $(x,l)^{\sigma\pi_{2}}\equiv(y,l+1)^{\sigma\pi_{2}}+1\pmod{n}$ or $(x,l)^{\sigma\pi_{2}}\equiv(y,l-1)^{\sigma\pi_{2}}-1\pmod{n}$, that is, $k_{3}\equiv k_{4}-1\pmod{n}$ or $k_{3}\equiv k_{4}+1\pmod{n}$. Since $v$ is adjacent to $x$ and $y$, $(v,l)$ is adjacent to $(x,l-1)$ and $(y,l-1)$, which implies that $(v,l)^{\sigma\pi_{2}}$ is adjacent to $(x,l-1)^{\sigma\pi_{2}}$ and $(y,l-1)^{\sigma\pi_{2}}$. Hence $(v,l)^{\sigma\pi_{2}}\equiv(x,l-1)^{\sigma\pi_{2}}+1\pmod{n}$ or $(v,l)^{\sigma\pi_{2}}\equiv(x,l-1)^{\sigma\pi_{2}}-1\pmod{n}$, that is, $k_{1}\equiv k_{3}+1\pmod{n}$ or $k_{1}\equiv k_{3}-1\pmod{n}$. Similarly, we have $k_{1}\equiv k_{4}+1\pmod{n}$ or $k_{1}\equiv k_{4}-1\pmod{n}$.
Since $w$ is adjacent to $x$ and $y$, $(w,l)$ is adjacent to $(x,l-1)$ and $(y,l-1)$, which implies that $(w,l)^{\sigma\pi_{2}}$ is adjacent to $(x,l-1)^{\sigma\pi_{2}}$ and $(y,l-1)^{\sigma\pi_{2}}$. Thus $(w,l)^{\sigma\pi_{2}}\equiv(x,l-1)^{\sigma\pi_{2}}+1\pmod{n}$ or $(w,l)^{\sigma\pi_{2}}\equiv(x,l-1)^{\sigma\pi_{2}}-1\pmod{n}$, that is, $k_{2}\equiv k_{3}+1\pmod{n}$ or $k_{2}\equiv k_{3}-1\pmod{n}$. Similarly, we have $k_{2}\equiv k_{4}+1\pmod{n}$ or $k_{2}\equiv k_{4}-1\pmod{n}$.
The two paragraphs above show that we have $k_{1}\equiv k_{2}\pmod{n}$, $k_{1}\equiv k_{2}-2\pmod{n}$, or $k_{1}\equiv k_{2}+2\pmod{n}$. If $k_{1}\equiv k_{2}\pmod{n}$, then we may assume $k_{1}\equiv k_{2}\equiv k_{3}-1\pmod{n}$ and $k_{1}\equiv k_{2}\equiv k_{4}-1\pmod{n}$, which contradicts the fact that $k_{3}\equiv k_{4}+1\pmod{n}$ or $k_{3}\equiv k_{4}-1\pmod{n}$. Hence $k_{1}\not\equiv k_{2}\pmod{n}$. Suppose that $k_{1}\equiv k_{2}-2\pmod{n}$. Then by $k_{1}\equiv k_{3}+1\pmod{n}$ or $k_{3}-1\pmod{n}$ and $k_{1}\equiv k_{4}+1\pmod{n}$ or $k_{4}-1\pmod{n}$, we obtain that $k_{2}\equiv k_{3}+1\pmod{n}$ or $k_{3}-3\pmod{n}$ and $k_{2}\equiv k_{4}+1\pmod{n}$ or $k_{4}-3\pmod{n}$. However, we have $k_{2}\equiv k_{3}-1\pmod{n}$ or $k_{3}+1\pmod{n}$ and $k_{2}\equiv k_{4}-1\pmod{n}$ or $k_{4}+1\pmod{n}$. Hence $k_{2}\equiv k_{3}+1\pmod{n}$ and $k_{2}\equiv k_{4}+1\pmod{n}$. Thus, $k_{3}\equiv k_{4}\pmod{n}$, which is a contradiction. Therefore, $k_{1}\not\equiv k_{2}-2$. Similarly, we can prove that $k_{1}\not\equiv k_{2}+2\pmod{n}$. This final contradiction shows that for any $(u,i)\in V(\Gamma\times C_{n})$ and $\sigma\in\mathrm{Aut}(\Gamma\times C_{n})$ we have $(u,i)^{\sigma\pi_{2}}\neq(u,i+2)^{\sigma\pi_{2}}$. Using this and Lemma 3.3, we obtain $\mathrm{Aut}(\Gamma\times C_{n})=\mathrm{P}(\Gamma,C_{n})$ immediately. ∎
Finally, Theorem 3.1 follows from Lemmas 3.2 and 3.4 immediately.
4. More results on the stability of $R$-thin connected graphs versus cycles
In this section we prove more results on the stability of $R$-thin connected graphs against cycles. The main results will be given in Theorems 4.3, 4.5 and 4.6.
In [14], Wilson introduced the concept of expected automorphisms of $\Gamma\times K_{2}$. We now generalize this concept from $\Gamma\times K_{2}$ to $\Gamma\times C_{n}$. Set $V(C_{n})=\{0,1,\ldots,n-1\}$. Recall that $\mathrm{Aut}\,(C_{n})$ is isomorphic to the dihedral group $D_{2n}$. For any $\delta\in D_{2n}$, let $\overline{\delta}$ be the permutation of $V(\Gamma\times C_{n})$ defined by
$(u,i)^{\overline{\delta}}=(u,i^{\delta}),\text{ for }(u,i)\in V(\Gamma\times C_{n}).$
It is readily seen that $\overline{\delta}$ is an automorphism of $\Gamma\times C_{n}$. Set $\overline{\mathrm{Aut}\,(C_{n})}=\{\overline{\delta}\,:\,\delta\in D_{2n}\}$. Then $\overline{\mathrm{Aut}\,(C_{n})}$ is a subgroup of $\mathrm{Aut}\,(\Gamma\times C_{n})$ isomorphic to $\mathrm{Aut}\,(C_{n})$. Similarly, for any $\sigma\in\mathrm{Aut}\,(\Gamma)$, let $\overline{\sigma}$ be the automorphism of $\Gamma\times C_{n}$ defined by
$(u,i)^{\overline{\sigma}}=(u^{\sigma},i),\text{ for }(u,i)\in V(\Gamma\times C_{n}).$
Set $\overline{\mathrm{Aut}\,(\Gamma)}=\{\overline{\sigma}\,:\,\sigma\in\mathrm{Aut}\,(\Gamma)\}$. Then $\overline{\mathrm{Aut}\,(\Gamma)}$ is a subgroup of $\mathrm{Aut}\,(\Gamma\times C_{n})$ isomorphic to $\mathrm{Aut}\,(\Gamma)$. Note that $\overline{\sigma}\ \overline{\delta}=\overline{\delta}\ \overline{\sigma}$.
It is known that $\mathrm{Aut}\,(\Gamma)\times\mathrm{Aut}\,(C_{n})$ is isomorphic to a subgroup of $\mathrm{Aut}\,(\Gamma\times C_{n})$. In fact, it is isomorphic to the subgroup
$\mathrm{R}(\Gamma,C_{n}):=\langle\overline{\mathrm{Aut}\,(\Gamma)},\overline{\mathrm{Aut}\,(C_{n})}\rangle$
of $\mathrm{Aut}\,(\Gamma\times C_{n})$ generalized by $\overline{\mathrm{Aut}\,(\Gamma)}$ and $\overline{\mathrm{Aut}\,(C_{n})}$. Note that $\mathrm{R}(\Gamma,C_{n})$ is a subgroup of $\mathrm{P}(\Gamma,C_{n})$ and is the normalizer of $\overline{\mathrm{Aut}\,(C_{n})}$ in $\mathrm{Aut}(\Gamma\times C_{n})$. The elements of $\mathrm{R}(\Gamma,C_{n})$ are called the expected automorphisms of $\Gamma\times C_{n}$, and the elements of $\mathrm{P}(\Gamma,C_{n})\setminus\mathrm{R}(\Gamma,C_{n})$ are called the unexpected automorphisms of $\Gamma\times C_{n}$. It is readily seen that for any $\gamma\in\mathrm{R}(\Gamma,C_{n})$, $u\in V(\Gamma)$, and $i,j\in V(C_{n})$, we have $(u,i)^{\gamma\pi_{1}}=(u,j)^{\gamma\pi_{1}}$. Of course, if $\mathrm{Aut}(\Gamma\times C_{n})=\mathrm{R}(\Gamma,C_{n})$, then $(\Gamma,C_{n})$ is stable; otherwise, $(\Gamma,C_{n})$ is unstable.
For a vertex $u$ of $\Gamma$ and a subgroup $H$ of $\mathrm{Aut}\,(\Gamma)$, let $u^{H}=\{u^{h}:h\in H\}$ be the $H$-orbit on $V(\Gamma)$ containing $u$. Let $\Gamma/H$ be the quotient graph of $\Gamma$ with respect to the partition $\{u^{H}:u\in V(\Gamma)\}$ of $V(\Gamma)$ ([14]). That is, $\Gamma/H$ is the graph with vertex set $\{u^{H}:u\in V(\Gamma)\}$ in which $\{u^{H},v^{H}\}$ is an edge if and only if $u^{H}\neq v^{H}$ and there exist $u^{\prime}\in u^{H}$ and $v^{\prime}\in v^{H}$ such that $\{u^{\prime},v^{\prime}\}\in E(\Gamma)$. In particular, for $\gamma\in\mathrm{Aut}\,(\Gamma)$, we write $\Gamma/\gamma$ in place of $\Gamma/\langle\gamma\rangle$. An arc of $\Gamma$ is an ordered pair of adjacent vertices $(u,v)$ of $\Gamma$. Denote by $A(\Gamma)$ the set of arcs of $\Gamma$. The following concepts are all extracted from [14].
Definition 4.1.
([14]) Let $\Gamma$ be a connected graph.
  1. (a)
    Let $\gamma\in\mathrm{Aut}\,(\Gamma)$ be an involution and $\Gamma_{1}=\Gamma/\gamma$. Let $\alpha_{1}\in\mathrm{Aut}\,(\Gamma_{1})$ and let $\pi$ be the natural projection from $\Gamma$ to $\Gamma/\gamma$ (that is, $\pi$ sends $u\in V(\Gamma)$ to $u^{\langle\gamma\rangle}\in V(\Gamma/\gamma)$). A covering permutation of $\alpha_{1}$ is a permutation $\alpha$ of $V(\Gamma)$ such that $(u^{\alpha})^{\pi}=(u^{\pi})^{\alpha_{1}}$ for every $u\in V(\Gamma)$. An anti-automorphism of $\Gamma$ is a permutation $\alpha$ of $V(\Gamma)$ which commutes with $\gamma$ such that $\{u^{\alpha},v^{\gamma\alpha}\}\in E(\Gamma)$ whenever $\{u,v\}\in E(\Gamma)$.
  2. (b)
    Let $\phi$ be an automorphism of $\Gamma$ with order at least $3$. If $\{u^{\phi},v^{\phi^{-1}}\}\in E(\Gamma)$ whenever $\{u,v\}\in E(\Gamma)$, then $\Gamma$ is called a (generalized) cross-cover of $\Gamma/\phi$.
  3. (c)
    Let $\Gamma_{1}$ be a graph and $\alpha\in\mathrm{Aut}\,(\Gamma_{1})$. Let $H=\mathbb{Z}^{n}_{2}$ and $\gamma\in\mathrm{Aut}\,(H)$. Let $L$ be a map from $V(\Gamma_{1})$ to the set of subgroups of $H$ such that $L(v^{\alpha})=L(v)^{\gamma}$ for all $v\in V(\Gamma_{1})$. Let $\omega$ be a map from $A(\Gamma_{1})$ to the set of subsets of $H$ such that $\omega(v,u)=\{x^{-1}:x\in\omega(u,v)\}$ for any $(u,v)\in A(\Gamma_{1})$ and there is a fixed element $h_{0}\in H\setminus(\cap_{v\in V}L(v))$ satisfying $\omega((u,v)^{\alpha})=\omega(u,v)^{\gamma}+h_{0}$ for all $(u,v)\in A(\Gamma_{1})$. Define $\Gamma=GV(\Gamma_{1},H,L,\omega)$ to be the graph with vertex set $V(\Gamma)=\{(v,L(v)+h):v\in V(\Gamma_{1}),h\in H\}$ and arc set
    $A(\Gamma)=\{((u,L(u)+h),(v,L(v)+h+x)):(u,v)\in A(\Gamma_{1}),h\in H,x\in\omega(u,v)\}.$
    We call $\Gamma=GV(\Gamma_{1},H,L,\omega)$ a twist of $\Gamma_{1}$.
The following lemma is extracted from [4] and [14].
Lemma 4.2.
The following hold:
  1. (a)
    If a graph has an anti-automorphism, or is a cross-cover or twist of some graph, then it must be unstable ([14, Theorems 2–4]).
  2. (b)
    If a graph is nontrivially unstable, then it has an anti-automorphism, or is a cross-cover of a twist of a graph ([14, Theorem 5]).
  3. (c)
    If $\Gamma$ is an $R$-thin connected non-bipartite graph, then the following conditions are equivalent:
    1. (i)
      $\Gamma$ is unstable;
    2. (ii)
      $\Gamma$ is nontrivially unstable;
    3. (iii)
      $\Gamma$ has a nondiagonal two-fold automorphism;
    4. (iv)
      $\Gamma$ has an anti-automorphism, or is a cross-cover or twist of some graph.
Proof.
A nontrivially unstable graph is precisely an $R$-thin connected non-bipartite unstable graph. Thus an $R$-thin connected non-bipartite graph is nontrivially unstable if and only if it is unstable. By [4, Theorem 3.2], a graph is unstable if and only if it has a nondiagonal two-fold automorphism. Combining all these with the statements in (a) and (b), we conclude that conditions (i)–(iv) in (c) are equivalent for $R$-thin connected non-bipartite graphs. ∎
Theorem 4.3.
Let $\Gamma$ be an $R$-thin connected non-bipartite graph. If $(\Gamma,K_{2})$ is unstable, then $(\Gamma,C_{2k})$ is unstable for every $k\geq 2$.
Proof.
Set $V(C_{2k})=\{0,1,\ldots,2k-1\}$. Suppose that $(\Gamma,K_{2})$ is unstable. Since $C_{4}$ is not $R$-thin, $(\Gamma,C_{4})$ is unstable, and hence the result is true when $k=2$. It remains to prove that $(\Gamma,C_{2k})$ is unstable for all $k\geq 3$. Since $(\Gamma,K_{2})$ is unstable, there is a nondiagonal two-fold automorphism $(\alpha_{1},\alpha_{2})$ of $\Gamma$. Define a permutation $\sigma$ of $V(\Gamma\times C_{2k})$ as follows: For any $(u,i)\in V(\Gamma\times C_{2k})$, if $i$ between 00 and $12k-$ 1 is odd, then set $(u,i)^{\sigma}=(u^{\alpha_{2}},i)$; if $i$ between 00 and $12k-$ 1 is even, then set $(u,i)^{\sigma}=(u^{\alpha_{1}},i)$. Since $\alpha_{1}$ and $\alpha_{2}$ are permutations of $V(\Gamma)$, $\sigma$ is indeed a permutation of $V(\Gamma\times C_{2k})$.
In order to prove $\sigma\in\mathrm{P}(\Gamma,C_{2k})$, we first prove $\sigma\in\mathrm{Aut}\,(\Gamma\times C_{2k})$. For any $\{(u,i),(v,i+1)\}\in E(\Gamma\times C_{2k})$, $\{(u,i)^{\sigma},(v,i+1)^{\sigma}\}$ is equal to $\{(u,i)^{\alpha_{1}},(v,i+1)^{\alpha_{2}}\}=\{(u^{\alpha_{1}},i),(v^{\alpha_{2}},i+1)\}$ or $\{(u,i)^{\alpha_{2}},(v,i+1)^{\alpha_{1}}\}=\{(u^{\alpha_{2}},i),(v^{\alpha_{1}},i+1)\}$. Thus $\{(u,i),(v,i+1)\}^{\sigma}\in E(\Gamma\times C_{2k})$ by the definition of $(\alpha_{1},\alpha_{2})$. Hence $\sigma$ maps edges of $\Gamma\times C_{2k}$ to edges of $\Gamma\times C_{2k}$. Moreover, since $\sigma$ is a permutation of $V(\Gamma\times C_{2k})$, it maps different edges of $\Gamma\times C_{2k}$ to different edges of $\Gamma\times C_{2k}$. Thus, $\sigma\in\mathrm{Aut}\,(\Gamma\times C_{2k})$. By the definition of $\sigma$, for any $(u,i)\in V(\Gamma\times C_{2k})$, we have $(u,i)^{\sigma\pi_{2}}=i$ and hence $\sigma\in\mathrm{P}(\Gamma,C_{2k})$. Moreover, since $(\alpha_{1},\alpha_{2})$ is nondiagonal, we have $\alpha_{1}\neq\alpha_{2}$ and so there exists $v\in V(\Gamma)$ such that $v^{\alpha_{1}}\neq v^{\alpha_{2}}$. This together with the definition of $\sigma$ implies that $(v,2i)^{\sigma\pi_{1}}\neq(v,2i+1)^{\sigma\pi_{1}}$ for $0\leq i\leq k$. Let $\delta$ be the unique automorphism of $C_{2k}$ satisfying $i^{\delta}\equiv i+1\pmod{2k}$ for any $i\in V(C_{2k})$. Then $\overline{\delta}\in\overline{\mathrm{Aut}\,(C_{2k})}$ is an automorphism of $\Gamma\times C_{2k}$ satisfying $(u,i)^{\bar{\delta}}=(u,i+1)$ for any $(u,i)\in V(\Gamma\times C_{2k})$. We have $\sigma\bar{\delta}\neq\bar{\delta}\sigma$. Thus $\sigma$ is an unexpected automorphism of $\Gamma\times C_{2k}$. Hence $(\alpha_{1},\alpha_{2},\alpha_{1},\alpha_{2},\ldots,\alpha_{1},\alpha_{2})$ is a nondiagonal $C_{2k}$-automorphism of $\Gamma$. Therefore, by [10, Lemma 2.6], $\Gamma\times C_{2k}$ is unstable. ∎
Note that $\mathrm{P}(\Gamma,\Sigma)\cong\mathrm{Aut}_{\Sigma}\,(\Gamma)\rtimes\mathrm{Aut}\,(\Sigma)$ for any graph pair $(\Gamma,\Sigma)$ (see [10, Lemma 2.4]). So, for any $\sigma\in\mathrm{P}(\Gamma,\Sigma)$, there exist $(\alpha_{1},\ldots,\alpha_{n})\in\mathrm{Aut}\,_{\Sigma}(\Gamma)$ and $\delta\in\mathrm{Aut}\,(\Sigma)$ such that $\sigma$ corresponds to $((\alpha_{1},\ldots,\alpha_{n}),\delta)\in\mathrm{Aut}_{\Sigma}\,(\Gamma)\rtimes\mathrm{Aut}\,(\Sigma)$. In the case when $\Sigma=C_{n}$, $(\alpha_{1},\ldots,\alpha_{n})$ is uniquely determined by $\sigma$ and $(u,i)^{\sigma\pi_{1}}=u^{\alpha_{i}}$ holds for any $(u,i)\in V(\Gamma\times C_{n})$.
Lemma 4.4.
Let $\Gamma$ be an $R$-thin connected non-bipartite graph and $k\geq 3$ an integer. If $2k$ is compatible with $\Gamma$ and $(\Gamma,C_{2k})$ is nontrivially unstable, then $\Gamma$ has an anti-automorphism, or is a cross-cover or twist of some graph.
Proof.
Set $V(C_{2k})=\{1,\ldots,2k\}$. Since $\Gamma$ is an $R$-thin connected non-bipartite graph and $2k$ is compatible with $\Gamma$, by Lemma 3.4, we have $\mathrm{Aut}(\Gamma\times C_{2k})=\mathrm{P}(\Gamma,C_{2k})$. Since $(\Gamma,C_{2k})$ is nontrivially unstable, there are unexpected automorphisms of $\Gamma\times C_{2k}$. Consider an arbitrary unexpected automorphism $\sigma\in\mathrm{Aut}(\Gamma\times C_{2k})$. For each $i\in V(C_{2k})$, define $u^{\alpha_{i}}=(u,i)^{\sigma\pi_{1}}$ for $u\in V(\Gamma)$. Note that for any $\{u,v\}\in E(\Gamma)$ and $\{i,i+1\}\in E(C_{2k})$ we have $\{u^{\alpha_{i}},v^{\alpha_{j}}\}=\{(u,i)^{\sigma\pi_{1}},(v,i+1)^{\sigma\pi_{1}}\}$. Since $\{(u,i),(v,i+1)\}\in E(\Gamma\times C_{2k})$ and $\sigma\in\mathrm{Aut}\,(\Gamma\times C_{2k})$, we see that $\{(u,i)^{\sigma},(v,i+1)^{\sigma}\}\in\mathrm{Aut}\,(\Gamma\times C_{2k})$ and so $\{u^{\alpha_{i}},v^{\alpha_{i+1}}\}\in E(\Gamma)$. Hence $\{u,v\}\in E(\Gamma)$ if and only if $\{u^{\alpha_{i}},v^{\alpha_{j}}\}\in E(\Gamma)$. So $(\alpha_{1},\ldots,\alpha_{2k})$ is a $C_{2k}$-automorphism of $\Gamma$. Since $\sigma$ is an unexpected automorphism, there exist distinct vertices $(u,i),(u,j)\in V(\Gamma\times C_{2k})$ such that $(u,i)^{\sigma\pi_{1}}\neq(u,j)^{\sigma\pi_{1}}$, that is, $u^{\alpha_{i}}\neq u^{\alpha_{j}}$ and therefore $\alpha_{i}\neq\alpha_{j}$. Let $\delta$ be the unique automorphism of $C_{2k}$ satisfying $i^{\delta}=2k-i\pmod{2k}$ for $i\in V(C_{2k})$. By Lemma 2.1, $(\alpha_{1},\ldots,\alpha_{2k})^{\delta}=(\alpha_{1},\alpha_{2k},\alpha_{2k-1},\ldots,\alpha_{k+2},\alpha_{k+1},\alpha_{k},\ldots,\alpha_{2})$ is a $C_{2k}$-automorphism of $\Gamma$. Since $\mathrm{Aut}\,_{C_{2k}}(\Gamma)$ is an group, $(\alpha^{-1}_{1},\alpha^{-1}_{2k},\alpha^{-1}_{2k-1},\ldots,\alpha^{-1}_{k+2},\alpha^{-1}_{k+1},\ldots,\alpha^{-1}_{2})$ is a $C_{2k}$-automorphism of $\Gamma$. Hence $({\rm id},\alpha_{2}\alpha^{-1}_{2k},\ldots,{\rm id},\alpha_{k+2}\alpha^{-1}_{k},\ldots,\alpha_{2k}\alpha^{-1}_{2})$ is also a $C_{2k}$-automorphism of $\Gamma$. Since $\Gamma$ is $R$-thin, by Lemma 2.3, this $C_{2k}$-automorphism of $\Gamma$ must be diagonal. This implies that $\alpha_{2}=\alpha_{2k},\alpha_{3}=\alpha_{2k-1},\ldots,\alpha_{k+2}=\alpha_{k}$. Thus $(\alpha_{1},\ldots,\alpha_{2k})=(\alpha_{1},\alpha_{2},\alpha_{3},\ldots,\alpha_{k},\alpha_{k+1},\alpha_{k},\alpha_{k-1},\ldots,\alpha_{3},\alpha_{2})$. Let $\tau$ be the unique automorphism of $C_{2k}$ satisfying $i^{\tau}\equiv i+2\pmod{2k}$ for $i\in V(C_{2k})$. Then $(\alpha_{1},\alpha_{2},\alpha_{3},\ldots,\alpha_{k},\alpha_{k+1},\alpha_{k},\alpha_{k-1},\ldots,\alpha_{3},\alpha_{2})^{\tau}=(\alpha_{3},\alpha_{4},\ldots,\alpha_{k},\alpha_{k-1},\alpha_{k-2},\ldots,\alpha_{1},\alpha_{2})$, which, by Lemma 2.1, is a $C_{2k}$-automorphism of $\Gamma$. Hence $(\alpha_{3}^{-1},\alpha_{4}^{-1},\ldots,\alpha_{k}^{-1},\alpha_{k-1}^{-1},\alpha_{k-2}^{-1},\ldots,\alpha_{1}^{-1},\alpha_{2}^{-1})$ is a $C_{2k}$-automorphism of $\Gamma$ and so $(\alpha_{1}\alpha_{3}^{-1},\alpha_{2}\alpha_{4}^{-1},\ldots,{\rm id},\alpha_{k+1}\alpha_{k-1}^{-1},\alpha_{k}\alpha_{k-2}^{-1},\ldots,\alpha_{3}\alpha_{1}^{-1},{\rm id})$ is also a $C_{2k}$-automorphism of $\Gamma$. Since $\Gamma$ is $R$-thin, $\alpha_{1}=\alpha_{3}=\cdots=\alpha_{2k-1}$ and $\alpha_{2}=\alpha_{4}=\cdots=\alpha_{2k}$. So we have proved that for any unexpected automorphism of $\Gamma\times C_{2k}$ there corresponds a $C_{2k}$-automorphism of $\Gamma$ with the form $(\alpha_{1},\alpha_{2},\ldots,\alpha_{1},\alpha_{2})$. Note that $\alpha_{1}\neq\alpha_{2}$ and $\{u^{\alpha_{1}},v^{\alpha_{2}}\}\in E(\Gamma)$ for any $\{u,v\}\in E(\Gamma)$. Thus $(\alpha_{1},\alpha_{2})$ is a two-fold automorphism of $\Gamma$. Since $\alpha_{1}\neq\alpha_{2}$, this two-fold automorphism is nondiagonal and hence $\Gamma$ is unstable. Thus, by Lemma 4.2(b), $\Gamma$ has an anti-automorphism, or is a cross-cover or twist of some graph. ∎
Theorem 4.5.
Let $\Gamma$ be an $R$-thin connected non-bipartite graph.
  1. (a)
    If $k\geq 3$ is an integer such that $2k$ is compatible with $\Gamma$, then $(\Gamma,C_{2k})$ is nontrivially unstable if and only if $(\Gamma,K_{2})$ is unstable.
  2. (b)
    If there exists an integer $k\geq 3$ such that $2k$ is compatible with $\Gamma$ and $(\Gamma,C_{2k})$ is nontrivially unstable, then $(\Gamma,C_{2l})$ is unstable for every $l\geq 3$.
Proof.
(a) Suppose $2k\geq 6$ is compatible with $\Gamma$. If $(\Gamma,C_{2k})$ is nontrivially unstable, then by Lemma 4.4, $\Gamma$ satisfies condition (iv) in Lemma 4.2(c). Hence, by Lemma 4.2(c), $\Gamma$ is unstable, that is, $(\Gamma,K_{2})$ is unstable. Conversely, if $(\Gamma,K_{2})$ is unstable, then by Theorem 4.3, $(\Gamma,C_{2k})$ is unstable. Since $2k$ is compatible with $\Gamma$, $\Gamma$ and $C_{2k}$ must be coprime. Note that $C_{2k}$ is $R$-thin as $2k\geq 6$ and $\Gamma$ is non-bipartite by our assumption. Hence $(\Gamma,C_{2k})$ is nontrivially unstable.
(b) Suppose $k\geq 3$ is an integer such that $2k$ is compatible with $\Gamma$ and $(\Gamma,C_{2k})$ is nontrivially unstable. Then, by (a), $(\Gamma,K_{2})$ is unstable. Hence, by Theorem 4.3, $(\Gamma,C_{2l})$ is unstable for every $l\geq 3$. ∎
Theorem 4.6.
Let $\Gamma$ be an $R$-thin connected graph and $k\geq 1$ an integer. If $12k+$ 1 is compatible with $\Gamma$, then $(\Gamma,C_{2k+1})$ is stable.
Proof.
Set $V(C_{2k+1})=\{0,1,\ldots,2k\}$. Since $\Gamma$ is an $R$-thin connected graph and $C_{2k+1}$ is an odd cycle, both $\Gamma$ and $C_{2k+1}$ are $R$-thin and at least one of them is non-bipartite. Since $12k+$ 1 is compatible with $\Gamma$, we have $\mathrm{Aut}(\Gamma\times C_{2k+1})=\mathrm{P}(\Gamma,C_{2k+1})$ by Lemma 3.4(a). Suppose for a contradiction that $(\Gamma,C_{2k+1})$ is unstable. Then $(\Gamma,C_{2k+1})$ is nontrivially unstable. So there exists an unexpected automorphism $\sigma$ of $\Gamma\times C_{2k+1}$. Define $u^{\alpha_{i}}=(u,i)^{\sigma\pi_{1}}$, $u\in V(\Gamma)$, for each $i\in V(C_{2k+1})$. Note that for any $\{u,v\}\in E(\Gamma)$ and $\{i,i+1\}\in E(C_{2k+1})$, we have $\{u^{\alpha_{i}},v^{\alpha_{i+1}}\}=\{(u,i)^{\sigma\pi_{1}},(v,i+1)^{\sigma\pi_{1}}\}$. Since $\sigma\in\mathrm{Aut}\,(\Gamma\times C_{2k+1})$ and $\{(u,i),(v,i+1)\}\in E(\Gamma\times C_{2k+1})$, we obtain that $\{u,v\}\in E(\Gamma)$ if and only if $\{(u,i)^{\sigma\pi_{1}},(v,i+1)^{\sigma\pi_{1}}\}\in E(\Gamma)$. So $(\alpha_{1},\alpha_{2},\ldots,\alpha_{2k+1})$ is a $C_{2k+1}$-automorphism of $\Gamma$. Since $\sigma$ is an unexpected automorphism of $\Gamma\times C_{2k+1}$, there exists an automorphism $\bar{\delta}\in\overline{\mathrm{Aut}\,(C_{2k+1})}$ such that $\sigma\bar{\delta}\neq\bar{\delta}\sigma$ and hence $\bar{\delta}^{-1}\sigma\bar{\delta}\neq\sigma$. Suppose that $(u,i)^{\bar{\delta}}=(u,j)$ for some $(u,i),(u,j)\in V(\Gamma\times C_{2k+1})$. Then $(u,i)^{\sigma\bar{\delta}}=((u,i)^{\sigma\pi_{1}},i)^{\bar{\delta}}=((u,i)^{\sigma\pi_{1}},j)\neq(u,i)^{\bar{\delta}\sigma}=(u,j)^{\sigma}$. Thus there exist distinct vertices $(u,i),(u,j)\in V(\Gamma\times C_{2k+1})$ such that $(u,i)^{\sigma\pi_{1}}\neq(u,j)^{\sigma\pi_{1}}$ and hence $\alpha_{i}\neq\alpha_{j}$. Hence $(\alpha_{1},\ldots,\alpha_{2k+1})$ is nondiagonal. Let $\tau$ be the permutation of $V(C_{2k+1})$ such that $i^{\tau}=2k+1-i\pmod{2k+1}$ for any $i\in V(C_{2k+1})$. Then $\{i^{\tau},j^{\tau}\}=\{2k+1-i,2k+1-j\}$ for any $\{i,j\}\in E(C_{2k+1})$. Since $\{i,j\}\in E(C_{2k+1})$, we have $j\equiv i+1\pmod{2k+1}$ or $j\equiv i-1\pmod{2k+1}$. Without loss of generality, we may assume $j\equiv i+1\pmod{2k+1}$. We can easily verify that $\{2k+1-i,2k+1-j\}=\{2k+1-i,2k-i\}$ is also an edge of $C_{2k+1}$. Since $\tau$ is a permutation of $V(C_{2k+1})$ and maps edges of $C_{2k+1}$ to edges of $C_{2k+1}$, we see that $\tau$ is an automorphism of $C_{2k+1}$. Since $\tau$ maps $i$ to $2k+1-i\pmod{2k+1}$ for each $i\in V(C_{2k+1})$, we have $(\alpha_{1},\ldots,\alpha_{2k+1})^{\tau}=(\alpha_{1},\alpha_{2k+1},\ldots,\alpha_{k+2},\alpha_{k+1},\ldots,\alpha_{2})$. Since $(\alpha_{1},\ldots,\alpha_{2k+1})\in\mathrm{Aut}\,_{C_{2k+1}}(\Gamma)$, it follows from Lemma 2.1 that $(\alpha_{1},\alpha_{2k+1},\ldots,\alpha_{k+2},\alpha_{k+1},\ldots,\alpha_{2})$ is a $C_{2k+1}$-automorphism of $\Gamma$. Moreover, since $\mathrm{Aut}\,_{C_{2k+1}}(\Gamma)$ is a group, $(\alpha^{-1}_{1},\alpha^{-1}_{2k+1},\ldots,\alpha^{-1}_{k+2},\alpha^{-1}_{k+1},\ldots,\alpha^{-1}_{2})$ is also a $C_{2k+1}$-automorphism of $\Gamma$. So $({\rm id},\alpha_{2}\alpha^{-1}_{2k+1},\ldots,\alpha_{k+1}\alpha^{-1}_{k+2},\alpha_{k+2}\alpha^{-1}_{k+1},\ldots,\alpha_{2k+1}\alpha^{-1}_{2})$ is a $C_{2k+1}$-automorphism of $\Gamma$. Since $\Gamma$ is $R$-thin, by Lemma 2.3, this $C_{2k+1}$-automorphism of $\Gamma$ must be diagonal. That is, $\alpha_{i}=\alpha_{2k+3-i}$ for $12\leq i\leq 2k+$ 1. So $(\alpha_{1},\ldots,\alpha_{2k+1})=(\alpha_{1},\alpha_{2},\ldots,\alpha_{k+1},\alpha_{k+1},\ldots,\alpha_{2})$. Obviously, the permutation $\phi$ of $V(C_{2k+1})$ which maps $i$ to $i-1\pmod{2k+1}$ for each $i\in V(C_{2k+1})$ is an automorphism of $C_{2k+1}$. By Lemma 2.1 again, it follows that $(\alpha_{1},\alpha_{2},\ldots,\alpha_{k+1},\alpha_{k+1},\ldots,\alpha_{2})^{\phi}=(\alpha_{2},\alpha_{3},\ldots,\alpha_{k+1},\alpha_{k+1},\ldots,\alpha_{2},\alpha_{1})$ is a $C_{2k+1}$-automorphism of $\Gamma$. Since $\mathrm{Aut}\,_{C_{2k+1}}(\Gamma)$ is a group, $(\alpha^{-1}_{2},\alpha^{-1}_{3},\ldots,\alpha^{-1}_{k+1},\alpha^{-1}_{k},\ldots,\alpha^{-1}_{1})$ is also a $C_{2k+1}$-automorphism of $\Gamma$. Hence
$(\alpha_{1}\alpha^{-1}_{2},\alpha_{2}\alpha^{-1}_{3},\ldots,{\rm id},\alpha_{k+1}\alpha^{-1}_{k},\alpha_{2}\alpha^{-1}_{2})$
is a $C_{2k+1}$-automorphism of $\Gamma$. Again, by Lemma 2.3, this $C_{2k+1}$-automorphism of $\Gamma$ must be diagonal. That is, $\alpha_{1}=\alpha_{2}=\cdots=\alpha_{2k+1}$ and so $(\alpha_{1},\ldots,\alpha_{2k+1})$ is diagonal, a contradiction. Combining this with Lemma 3.2(a), we conclude that $(\Gamma,C_{2k+1})$ is stable. ∎
Acknowledgement   This work was funded in part by the National Natural Science Foundation of China (Grants No.12071194, 11571155).